The Chow Ring of a Classifying Space

Burt Totaro

For any linear algebraic group G, we define a ring CH^{∗}BG, the ring of characteristic classes with values in the Chow ring (that is, the ring of algebraic cycles modulo rational equivalence) for principal G-bundles over smooth algebraic varieties. We show that this coincides with the Chow ring of any quotient variety (V − S)/G in a suitable range of dimensions, where V is a representation of G and S is a closed subset such that G acts freely outside S. As a result, computing the Chow ring of BG amounts to the computation of Chow groups for a natural class of algebraic varieties with a lot of torsion in their cohomology. Almost nothing is known about this in general. For G an algebraic group over the complex numbers, there is an obvious ring homomorphism CH^{∗}BG → H^{∗}(BG, Z). Less obviously, using the results of [42], this homomorphism factors through the quotient of the complex cobordism ring MU^{∗}BG by the ideal generated by the elements of negative degree in the coefficient ring MU^{∗} = Z[x_{1}, x_{2}, …], that is, through the ring MU^{∗}BG⊗_{MU^{∗}}Z. (For clarity, let us mention that the classifying space of a complex algebraic group is homotopy equivalent to that of its maximal compact subgroup. Moreover, every compact Lie group arises as the maximal compact subgroup of a unique complex reductive group.)

The most interesting result of this paper is that in all the examples where we can compute the Chow ring of BG, it maps isomorphically to the topologically defined ring MU^{∗}BG ⊗_{MU^{∗}} Z. Namely, this is true for finite abelian groups, the symmetric groups, tori, GL(n, C), Sp(2n, C), O(n), SO(2n + 1), and SO(4). (The computation of the Chow ring for SO(2n + 1) is the result of discussion between me and Rahul Pandharipande. Pandharipande then proceeded to compute the Chow ring of SO(4), which is noticeably more difficult [31].) We also get various additional information about the Chow rings of the symmetric groups, the group G_{2}, and so on.

Unfortunately, the map CH^{∗}BG → MU^{∗}BG ⊗_{MU^{∗}} Z is probably not always an isomorphism, since this would imply in particular that MU^{∗}BG is concentrated in even degrees. By Ravenel, Wilson, and Yagita, MU^{∗}BG is concentrated in even degrees if all the Morava K-theories of BG are concentrated in even degrees [33]. The latter statement, for all compact Lie groups G, was a plausible conjecture of Hopkins, Kuhn, and Ravenel [19], generalizing the theorem of Atiyah-Hirzebruch-Segal on the topological K-theory of BG [1], [5]. But it has now been disproved by Kriz [22], using the group G of strictly upper triangular 4 × 4 matrices over Z/3.

1991 Mathematics Subject Classification. Primary 14C25; Secondary 14L30, 20J06, 55N22.

The author was supported in part by a Sloan Research Fellowship.

Nonetheless, there are some reasonable conjectures to make. First, if G is a complex algebraic group such that the complex cobordism ring of G is concentrated in even degrees, say after tensoring with Z_{(p)} (Z localized at p) for a fixed prime number p, then the homomorphism

CH^{∗}BG → MU^{∗}BG ⊗_{MU^{∗}} Z

should become an isomorphism after tensoring with Z_{(p)}. Second, we can hope that the Chow ring of BG for any complex algebraic group G has the good properties which complex cobordism was formerly expected to have. In particular, the Chow ring of BG for a finite group G should be additively generated by transfers to G of Chern classes of representations of subgroups of G (see section 4). This conjecture suggests thinking of the Chow ring of BG as a better-behaved substitute for the ordinary group cohomology ring.

In many cases where we compute the Chow ring of BG, it maps injectively to the integral cohomology of BG: that is in fact true for all the examples mentioned above (finite abelian groups, the symmetric groups, tori, GL(n, C), Sp(2n, C), O(n), SO(2n + 1), and SO(4)). In such cases, we can say that the Chow ring of BG is the image of MU^{∗}BG in H^{∗}(BG, Z). Nonetheless, that is not true in general: for G = Z/2 × SO(4), the Chow ring of BG is equal to MU^{∗}BG ⊗_{MU^{∗}} Z, but this does not inject into H^{∗}(BG, Z). This observation was used to give a topological construction of elements of the Griffiths group of certain smooth projective varieties in [42].

The motivation for looking at these questions is the problem of determining which torsion cohomology classes on a smooth projective variety X can be represented by algebraic cycles. In particular, the image of CH^{1}X → H^{2}(X, Z) always contains the torsion subgroup of H^{2}(X, Z), but Atiyah and Hirzebruch [4] showed that for i ≥ 2 there are varieties X and torsion elements of H^{2i}(X, Z) which are not in the image of CH^{i}X. Their examples are Godeaux-Serre varieties, that is, quotients of smooth complete intersections by free actions of finite groups G. This leads to the question of actually computing the Chow groups of Godeaux-Serre varieties to the extent possible. It turns out that this problem almost completely reduces to the problem of computing the Chow ring of BG if we believe a suitable version of a conjecture of Nori’s (section 5). For the codimension 2 Chow group of Godeaux-Serre varieties, we can prove a strong relation to the codimension 2 Chow group of BG.

The outline of the paper is as follows. In section 1, we define the Chow ring of BG for an algebraic group G over any field. In section 2, restricting to groups over the complex numbers, we construct the factorization CH^{∗}BG → MU^{∗}BG ⊗_{MU^{∗}} Z → H^{∗}(BG, Z). The geometric work has already been done in [42], so unfortunately what remains is a technical argument about inverse limits. Section 3 describes the analogous notion of the algebraic K-group K_{0} of BG, which has been completely computed by Merkurjev [26]. We use this to show, for example, that the image of the homomorphism CH^{i}BG → (MU^{∗}BG⊗_{MU^{∗}} Z)^{2i} contains (i − 1)! times the latter group. Moreover, the map is an isomorphism when i is 1 or 2. Section 4 discusses the conjecture that CH^{∗}BG is generated by transferred Euler classes for G finite. In section 5, we conjecture the relation of the Chow ring of BG to the Chow ring of Godeaux-Serre varieties, and we prove most of it in the case of CH^{2}. Section 6 discusses the Chow groups of products in the cases we need. Sections 7 to 11 lead up to the proof that CH^{∗}BS_{n} → MU^{∗}BS_{n} ⊗_{MU^{∗}} Z is an isomorphism, where S_{n} denotes the symmetric group. Section 12 gives a curious description of the ring CH^{∗}BS_{n} ⊗ Z/2, and section 13 analyzes the Chow ring of the symmetric group over base fields other than the complex numbers. Section 14 proves the last general result on the Chow ring of classifying spaces, an explicit upper bound for the degree of generators for this ring. Nothing similar is known for either ordinary cohomology or complex cobordism. The bound is used in sections 15 and 16 to compute CH^{∗}BG for most of the classical groups G.

The inspiration for the definition of the Chow ring of BG came from Bogomolov’s work on the problem of rationality for quotient varieties [6]. I thank Dan Edidin, Hélène Esnault, Bill Fulton, Bill Graham, Bruno Kahn, Nick Kuhn, Doug Ravenel, Frank Sottile, Steve Wilson, and Nobuaki Yagita for telling me about their work. I have included Chris Stretch’s unpublished proof that MU^{∗}BS_{6} is not generated by Chern classes in section 4.

Contents

1. Definition of the Chow ring of an algebraic group 3 2. The factorization CH^{∗}BG → MU^{∗}BG ⊗_{MU^{∗}} Z → H^{∗}(BG, Z) 7 3. K_{0}BG, and the Chow ring with small primes inverted 9 4. Transferred Euler classes 14 5. Chow groups of Godeaux-Serre varieties 15 6. Some examples of the Chow Künneth formula 17 7. Chow groups of cyclic products, introduction 19 8. Chow groups of cyclic products, concluded 22 9. Wreath products 23 10. The symmetric group: injectivity 24 11. The Chow ring of the symmetric group 24 12. The ring CH^{∗}(BS_{n}) ⊗ Z/2 25 13. The Chow ring of the symmetric group over a general field 26 14. Generators for the Chow ring of BG 28 15. The Chow rings of the classical groups 30 16. Other connected groups 31 References 32

1. Definition of the Chow ring of an algebraic group

Fix a field k. We will work throughout in the category of separated schemes of finite type over k, called “varieties” for short. (We will try to say explicitly when we use the word “variety” in the more standard sense of a reduced irreducible separated scheme of finite type over k.) For any variety X, let CH_{i}X denote the group of i-dimensional algebraic cycles on X modulo rational equivalence [13], called the ith Chow group of X. If X is smooth of dimension n, we define the ith Chow cohomology group of X to be CH^{i}X := CH_{n−i}X. Then CH^{∗}X is a graded ring, called the Chow ring of X. The Chow ring is contravariant for arbitrary maps of smooth varieties. Let G be a linear algebraic group over k. A principal G-bundle over an algebraic variety X is a variety E with free G-action such that X = E/G.

Theorem 1.1. Let G be a linear algebraic group over a field k. Let V be any representation of G over k such that G acts freely outside a G-invariant closed subset S ⊂ V of codimension ≥ s. Suppose that the geometric quotient (V − S)/G (in the sense of [28]) exists as a variety over k. Then the ring CH^{∗}(V − S)/G, restricted to degrees less than s, is independent (in a canonical way) of the representation V and the closed subset S.

Moreover, such pairs (V, S) exist with the codimension of S in V arbitrarily large (see Remark 1.4, below). So we can make the following definition:

Definition 1.2. For a linear algebraic group G over a field k, define CH^{i}BG to be the group CH^{i}(V − S)/G for any (V, S) as in Theorem 1.1 such that S has codimension greater than i in V.

Then CH^{∗}BG forms a ring by Theorem 1.1, called the Chow ring of BG. The ring CH^{∗}BG depends on the field k; when we want to indicate this explicitly in the notation, we will write CH^{∗}(BG)_{k}.

For reductive groups G, such as finite groups, we can provide a more convincing justification for the name: the Chow ring of BG is equal to the ring of characteristic classes for principal G-bundles over smooth quasi-projective varieties, in the following sense.

Theorem 1.3. Let G be a reductive group over a field k. Then the above group CH^{i}BG is naturally identified with the set of assignments α to every smooth quasiprojective variety X over k with a principal G-bundle E over X of an element α(E) ∈ CH^{i}X, such that for any map f : Y → X we have α(f^{∗}E) = f^{∗}(α(E)). The ring structure on CH^{∗}BG is the obvious one.

Remark 1.4. If G is finite, then the geometric quotient V/G exists as an affine variety for all representations V of G [28]. So (V −S)/G is a quasi-projective variety for all closed subsets S ⊂ V such that G acts freely on V − S.

For any linear algebraic group G and any positive integer s, there is a representation V of G and a closed subset S ⊂ V of codimension at least s such that G acts freely on V − S and (V − S)/G exists as a quasi-projective variety over k. For example, let W be any faithful representation of G, say of dimension n, and let V = Hom(A^{N+n}, W) ≅ W^{⊕N+n} for N large. Let S be the closed subset in V of non-surjective linear maps A^{N+n} → W. Then the codimension of S in V goes to infinity as N goes to infinity. Also, (V − S)/G exists as a quasi-projective variety. Indeed, we can view it as the homogeneous space GL(N + n)/(H × G) where H is the subgroup of GL(N + n) which acts as the identity on a given ndimensional quotient space of A^{N+n}, so that H is an extension of GL(N) by an additive group Hom(A^{N}, A^{n}). Here any quotient of a linear algebraic group by a closed subgroup exists as a quasi-projective variety by [43], pp. 121-122. Now the Chow ring of a variety pulls back isomorphically to the Chow ring of an affine bundle over it, as we see by viewing an affine bundle as the difference of two projective bundles. It follows that the Chow ring of (V − S)/G agrees with the Chow ring of GL(N + n)/(GL(N) × G) for this choice of V and S.

Remark 1.5. Edidin and Graham generalized the above definition of CH^{∗}BG, thought of as the G-equivariant Chow ring of a point. Namely, they defined the G-equivariant Chow ring of a smooth G-variety X by

CH_{G}^{i}X = CH^{i}(X × (V − S))/G

for any (V, S) as above such that the codimension of S in V is greater than i [11].

Proof. (Theorem 1.1) We have to show that in codimension i < s, the Chow groups of (V − S)/G are independent of the subset S and the representation V. The independence of S just uses the basic exact sequence for Chow groups:

CH_{∗}Y → CH_{∗}X → CH_{∗}U → 0.

Here X is a variety, Y a closed subset, and U = X−Y. Apply this to X = (V −S)/G and U = (V −S′)/G, where S′ is some larger G-invariant subset of codimension ≥ s; since CH_{∗}(S′ − S)/G vanishes in dimensions greater than dim (V − S)/G − s, the Chow groups of (V − S)/G map isomorphically to the Chow groups of (V − S′)/G in codimension less than s. So the Chow ring of (V − S)/G is independent of S in the range we consider.

The independence of V follows from the double fibration construction, used for example by Bogomolov [6]. That is, consider any two representations V and W of G such that G acts freely outside subsets S_{V} and S_{W} of codimension ≥ s and such that the quotients (V − S_{V})/G and (W − S_{W})/G exist as varieties. Then consider the direct sum V ⊕ W. The quotient variety ((V − S_{V}) × W)/G exists, being a vector bundle over (V −S_{V})/G, and likewise the quotient variety (V ×(W −S_{W}))/G exists, as a vector bundle over (W − S_{W})/G. Independence of S (applied to the representation V ⊕W) shows that these two vector bundles have the same Chow ring in degrees less than s. Since the Chow ring of a variety X pulls back isomorphically to the Chow ring of any vector bundle over X [13], (V − S_{V})/G and (W − S_{W})/G have the same Chow ring in degrees less than s. QED.

Proof. (Theorem 1.3) Let A^{∗}BG (for the duration of this proof) denote the ring of characteristic classes for principal G-bundles over quasi-projective varieties, as defined in the statement of Theorem 1.3. We know that the Chow ring of quotient varieties (V − S)/G, when they exist, is independent of V and S in degrees less than the codimension of S. We have a natural homomorphism

A^{∗}BG → CH^{∗}(V − S)/G

when the quotient variety (V − S)/G is quasi-projective. Since there are some quotients (V − S)/G with the codimension of S arbitrarily large which are quasiprojective varieties (Remark 1.4, above), we get a natural homomorphism

A^{∗}BG → CH^{∗}BG,

by the definition of the latter groups. We will show that this homomorphism is an isomorphism.

Lemma 1.6. Let G be a reductive group, and let s be a positive integer. For any quasi-projective variety X with a principal G-bundle E, there is an affine-space bundle π : X′ → X and a map f : X′ → (V − S)/G such that π^{∗}E ≅ f^{∗}F, where F is the obvious principal G-bundle over (V − S)/G. Here V is a representation V of G and S is a closed subset of codimension ≥ s such that G acts freely on V − S and the quotient variety (V − S)/G exists as a quasi-projective variety.

Proof. By Jouanolou’s trick [20], for any quasi-projective variety X there is an affine-space bundle π : X′ → X such that X′ is affine. Pick any such X′ associated to the given variety X. Then X′ is an affine variety with a principal G-bundle π^{∗}E; that is, we have X′ = Y′/G for an affine variety Y′ with free G-action.

Let Y′ ⊂ A^{n} be any embedding of Y′ in affine space, given by a finitedimensional linear space W_{0} ⊂ O(Y′). Let W ⊂ O(Y′) be the G-linear span of W_{0}; then W is a finite-dimensional (by [28], pp. 25-26) representation of G which gives a G-equivariant embedding Y′ ↪ W^{∗}. Since G acts freely on Y′, the orbits of G on points of Y′ are closed in Y′ and hence in W^{∗}. That is, Y′ lies in the G-stable subset of W^{∗} in the sense of geometric invariant theory, which applies since G is reductive [28]. The theory shows that the G-stable set is an open subset of W^{∗} on which G acts properly, so there is a smaller open subset W^{∗} − S ⊂ W^{∗} containing Y′ on which G acts freely. Moreover, the geometric quotient (W^{∗} − S)/G exists as a quasi-projective variety [28]. By adding another representation of G to W^{∗} if necessary, we can arrange that S has codimension at least any given number s. Thus we have a map Y = Y′/G → (V − S)/G with the desired properties. QED.

It follows immediately from the lemma that the homomorphism A^{∗}BG → CH^{∗}BG is injective. Namely, if α is a characteristic class in A^{i}BG which maps to 0 in CH^{i}BG, then by definition of the latter group, α gives 0 on the quasi-projective quotients (V − S)/G with S of codimension greater than i. By pulling back, the characteristic class gives 0 on all smooth affine varieties with principal G-bundles. Since the Chow ring of a smooth variety maps isomorphically to that of an affine bundle over it, the characteristic class gives 0 on arbitrary smooth quasi-projective varieties with principal G-bundles. By our definition of A^{∗}BG, this means that α = 0.

To prove that A^{i}BG → CH^{i}BG is surjective, we argue as follows. Given any element α ∈ CH^{i}BG, which gives us an element of CH^{i}(V − S)/G for all quotient varieties (V − S)/G with S of codimension greater than i, we want to produce an element of CH^{i}X for arbitrary smooth quasi-projective varieties X with G-bundles E. The lemma explains how to do this: choose an affine-space bundle π : X′ → X such that there is a G-equivariant embedding f : Y′ → V (where X′ = Y′/G) for some representation V of G which is free outside a subset S of codimension greater than i, and pull back α ∈ CH^{i}(V − S)/G to CH^{i}X′ ≅ CH^{i}X. The only problem is to show that this is well-defined. We first show that it is well-defined for a given affine Y′, independent of the choice of the embedding f. Indeed, given any two embeddings f : Y′ → V and g : Y′ → W, we can consider the diagonal embedding h : Y′ → V ⊕ W; the commutative diagram

CH^{i}((V ⊕ W) − S_{V ⊕W})/G  CH^{i}(V − S_{V})/G
≅↓  ↓
CH^{i}(W − S_{W})/G  ⟶  CH^{i}Y′/G

shows that the two embeddings give the same element α ∈ CH^{i}Y′/G. Then, at last, if we have two different affine bundles over the same quasi-projective variety X which are both affine varieties, there is a product affine bundle which is an affine bundle over each one of them and in particular is an affine variety. It follows that α ∈ CH^{i}X is well-defined for each smooth quasi-projective variety X with a principal G-bundle. That is, A^{∗}BG → CH^{∗}BG is surjective as well as injective. QED.

2. The factorization CH^{∗}BG → MU^{∗}BG ⊗_{MU^{∗}} Z → H^{∗}(BG, Z)

By [42], for every smooth complex algebraic variety X, the cycle map CH^{∗}X → H^{∗}(X, Z) factors naturally through a more refined topological invariant of X, the ring MU^{∗}X ⊗_{MU^{∗}} Z. I will briefly describe what this means before explaining how it extends to classifying spaces BG.

Namely, complex cobordism MU^{∗}X is a cohomology theory in the topological sense; it is a graded ring associated to any topological space X (see Stong [36] for a general reference). When X is a complex manifold of complex dimension n, there is a simple interpretation of the group MU^{i}X as the bordism group of real (2n − i)-manifolds M with a complex linear structure on the stable tangent bundle T M ⊕ R^{N} and with a continuous proper map M → X. Another important way to think of complex cobordism is as the universal complex-oriented cohomology theory, where a cohomology theory h^{∗} is complex-oriented if we are given an element c_{1} in the reduced cohomology h̃^{∗}(CP^{∞}) whose restriction to h̃^{∗}CP^{1} freely generates the h^{∗}-module h̃^{∗}(CP^{1}). The familiar cohomology theories, integral cohomology and topological K-theory, are complex-oriented in a natural way, which gives natural transformations MU^{∗}X → H^{∗}(X, Z) and MU^{∗}X → K^{∗}X. When X is a point, the ring MU^{∗} := MU^{∗}(point) is a polynomial ring over Z on infinitely many generators, one in each degree −2i for i ≥ 1 (corresponding to some smooth compact complex ifold, in terms of the bordism definition of MU^{∗}). The map MU^{∗} → H^{∗}(point, Z) = Z is the identity in degree 0 and (of course) 0 in degrees less than 0, whereas the map MU^{∗} → K^{∗} takes the class of a complex i-fold in MU^{−2i} to its Todd genus in K^{−2i} = Z. We will need the map to K-theory in section 3.

Using the trivial homomorphism MU^{∗} → Z coming from integral cohomology, we can form the tensor product ring MU^{∗}X ⊗_{MU^{∗}} Z; this is a quotient ring of MU^{∗}X which maps to integral cohomology H^{∗}(X, Z). The map is an isomorphism for all spaces X with torsion-free integral cohomology, but not in general (as we will see for many classifying spaces BG), although it always becomes an isomorphism after tensoring with the rationals. The main result of [42] is that, for any smooth complex algebraic variety X, the cycle map CH^{∗}X → H^{∗}(X, Z) factors naturally through the ring MU^{∗}X ⊗_{MU^{∗}} Z. The factorization maps each codimension i subvariety Z in X to the class in MU^{2i}X of any resolution of singularities for Z, using the above bordism interpretation of MU^{2i}X. This class depends on the choice of resolution, but its image in MU^{∗}X ⊗_{MU^{∗}} Z does not, and is invariant under rational (or algebraic) equivalence of cycles.

Applying this to the varieties (V − S)/G for representations V of a complex algebraic group G, we get a natural map CH^{∗}BG → \underleftarrow{lim} MU^{∗}(V − S)/G ⊗_{MU^{∗}} Z, the inverse limit being taken over the representations V of G. One can identify this inverse limit with MU^{∗}BG ⊗_{MU^{∗}} Z, giving the following theorem. Unfortunately, this identification of the inverse limit is quite technical to prove, and so this section might well be skipped on first reading.

Theorem 2.1. For any complex algebraic group G, the ring homomorphism CH^{∗}BG → H^{∗}(BG, Z) factors naturally through MU^{∗}BG ⊗_{MU^{∗}} Z.

Proof. Here MU^{∗}BG must be viewed as a topological abelian group using Landweber’s description of it as MU^{∗}BG = \underleftarrow{lim} MU^{∗}(BG)_{n}, where (BG)_{n} denotes the n-skeleton of a CW-complex representing BG (see Lemma 2.2, below). The tensor product MU^{∗}BG ⊗_{MU^{∗}} Z must be defined as a topological tensor product:

MU^{∗}BG ⊗_{MU^{∗}} Z := \underleftarrow{lim}(im(MU^{∗}BG → MU^{∗}(BG)_{n}) ⊗_{MU^{∗}} Z),

as in [42], p. 481. And what is needed to prove the lemma is to show that, for any compact Lie group G, the ring MU^{∗}BG ⊗_{MU^{∗}} Z maps isomorphically to the inverse limit

\underleftarrow{lim}(MU^{∗}(BG)_{n} ⊗_{MU^{∗}} Z).

Indeed, this last inverse limit is clearly the same as the inverse limit of the rings MU^{∗}(V − S)/G ⊗_{MU^{∗}} Z, which is the ring to which CH^{∗}BG evidently maps, as explained above. (It is convenient to casually identify a compact Lie group with its complexification here; this is justified since the two groups have homotopy equivalent classifying spaces.)

We have only a rather elaborate proof that MU^{∗}BG ⊗_{MU^{∗}} Z maps isomorphically to the above inverse limit. In fact, it seems that one can show that the ring MU^{∗}BG ⊗_{MU^{∗}} Z is a finitely generated abelian group in each degree, which means that these inverse systems are in some sense trivial. We will not need that result in this paper. More strongly, we can hope that MU^{∗}BG ⊗_{MU^{∗}} Z is a finitely generated Z-algebra with generators in positive degree.

We need Landweber’s general results on the complex cobordism of classifying spaces [24]:

Lemma 2.2. Let G be a compact Lie group, and let (BG)_{n} denote the nskeleton of a CW complex BG. Then

MU^{∗}BG = \underleftarrow{lim} MU^{∗}(BG)_{n}.

Moreover, for each positive integer n, there is an N ≥ n such that, in all degrees at once,

im(MU^{∗}BG → MU^{∗}(BG)_{n}) = im(MU^{∗}(BG)_{N} → MU^{∗}(BG)_{n}).

To prove Theorem 2.1, we need to show that the natural map

MU^{∗}BG ⊗_{MU^{∗}} Z → \underleftarrow{lim}(MU^{∗}(BG)_{n} ⊗_{MU^{∗}} Z)

is an isomorphism.

Injectivity is fairly easy. Suppose that x ∈ MU^{∗}BG ⊗_{MU^{∗}} Z maps to 0 in the above inverse limit. By definition, x itself is a compatible sequence of elements x_{n} ∈ im(MU^{∗}BG → MU^{∗}(BG)_{n}) ⊗_{MU^{∗}} Z. Lift x_{n} to an element of the same name in im(MU^{∗}BG → MU^{∗}(BG)_{n}). We are assuming that, for each n, x_{n} belongs to MU^{<0} · MU^{∗}(BG)_{n}, and we want to show that for each n, x_{n} belongs to MU^{<0} · im(MU^{∗}BG → MU^{∗}(BG)_{n}).

By Lemma 2.2, for each n there is an N ≥ n such that

im(MU^{∗}BG → MU^{∗}(BG)_{n}) = im(MU^{∗}(BG)_{N} → MU^{∗}(BG)_{n}),

in all degrees at once. By hypothesis, we know that x_{N} belongs to MU^{<0} · MU^{∗}(BG)_{N}, and we can restrict this equality to (BG)_{n}. We deduce the desired conclusion about x_{n}, using the compatibility between x_{N} and x_{n}. So the above map is injective.

Now we only need to prove surjectivity of the homomorphism

MU^{∗}BG ⊗_{MU^{∗}} Z → \underleftarrow{lim}(MU^{∗}(BG)_{n} ⊗_{MU^{∗}} Z).

Suppose we are given an element of the inverse limit on the right. Let x_{n} ∈ MU^{∗}(BG)_{n}, n ≥ 1, be a choice of element representing the given element of MU^{∗}(BG)_{n} ⊗_{MU^{∗}} Z. By Lemma 2.2 again, for each n there is an N(n) ≥ n such that

im(MU^{∗}BG → MU^{∗}(BG)_{n}) = im(MU^{∗}(BG)_{N(n)} → MU^{∗}(BG)_{n}),

in all degrees at once. Clearly we can assume that N(1) < N(2) < · · ·. Let y_{n} be the restriction of x_{N(n)} to MU^{∗}(BG)_{n}; by definition of N(n), y_{n} lifts to an element of MU^{∗}BG, which we also call y_{n}.

Moreover, we know a certain compatibility among the elements y_{n} ∈ MU^{∗}BG: all the elements y_{m} for m ≥ n have the same restriction to the ring MU^{∗}(BG)_{n}⊗_{MU^{∗}} Z, since the x’s form an inverse system in these tensor product rings. Equivalently, for any positive integers m ≥ n, we can write the restriction of y_{m} − y_{n} to MU^{∗}(BG)_{n} as a finite sum of products of elements of MU^{<0} with elements of MU^{∗}(BG)_{n}. For each positive integer n, apply this fact to the positive integers N(n + 1) ≥ N(n): we find that the restriction of y_{N(n+1)} − y_{N(n)} to MU^{∗}(BG)_{N(n)} belongs to MU^{<0} · MU^{∗}(BG)_{N(n)}, meaning by this the additive group generated by such products. Then the definition of N(n) implies that upon further restricting to MU^{∗}(BG)_{n}, the element y_{N(n+1)} − y_{N(n)} belongs to MU^{<0} · im(MU^{∗}BG → MU^{∗}(BG)_{n}).

Therefore, we can define one last sequence of elements of MU^{∗}BG, z_{1}, z_{2}, …, by z_{1} = y_{N(1)}, z_{2} equals y_{N(2)} minus some element of MU^{<0} · MU^{∗}BG such that z_{1} and z_{2} have the same restriction to (BG)_{1}, z_{3} equals y_{N(3)} minus some element of MU^{<0} · MU^{∗}BG such that z_{3} and z_{2} have the same restriction to (BG)_{2}, and so on. Thus the z_{i}’s form an element of \underleftarrow{lim} MU^{∗}(BG)_{n} = MU^{∗}BG (the equality by Lemma 2.2). The resulting element maps to the element of \underleftarrow{lim}(MU^{∗}(BG)_{n} ⊗_{MU^{∗}} Z) we started with. Theorem 2.1 is proved. QED.

3. K_{0}BG, and the Chow ring with small primes inverted

One can define a ring K_{0}BG analogous to CH^{∗}BG using the algebraic K-group K_{0} in place of the Chow ring. The ring K_{0}BG can be completely computed using a theorem of Merkurjev [26]. The result is an algebraic analogue of the theorem of Atiyah-Hirzebruch-Segal on the topological K-theory of classifying spaces [1], [5]. We will use the calculation of K_{0}BG to derive some good information on CH^{∗}BG and MU^{∗}BG ⊗_{MU^{∗}} Z with small primes inverted.

For any linear algebraic group G over a field k, we define

K_{0}BG = \underleftarrow{lim} K_{0}(V − S)/G,

where the inverse limit runs over representations V of G over k. Here K_{0}X is the Grothendieck group of algebraic vector bundles over a variety. This (standard) notation has the peculiarity that for a variety X over the complex numbers, there is natural map from the algebraic K-group K_{0}X to the topological K-group K^{0}X.

To state the calculation of K_{0}BG, for any algebraic group G over a field k, let R(G) denote the representation ring of G over k (the Grothendieck group of finitedimensional representations of G over k). Thus R(G) is the free abelian group on the set of irreducible representations of G over k. Let R(G)^{∧} be the completion of R(G) with respect to powers of the augmentation ideal ker(R(G) → Z).

Theorem 3.1. For any algebraic group G over a field k, there is an isomorphism

R(G)^{∧} → K_{0}BG.

For comparison, Atiyah, Hirzebruch, and Segal proved that for a complex algebraic group G (or, equivalently, a compact Lie group), the completed representation ring R(G)^{∧} maps isomorphically to the topological K-group K^{0}BG, whereas K^{1}BG = 0 [1], [5]. Thus Theorem 3.1 implies that for a complex algebraic group G, the algebraic K-group K_{0}BG maps isomorphically to the topological K-group K^{0}BG.

Proof. Merkurjev computed, in particular, the algebraic K-group K_{0} for any homogeneous space of the form GL(n)/G:

K_{0}(GL(n)/G) = Z ⊗_{R(GL(n))} R(G)

[26]. By Remark 1.4, applied to K_{0} rather than the Chow ring, we can compute K_{0}BG as

K_{0}BG = \underleftarrow{lim}_{N}K_{0}(GL(N + n)/GL(N) × G)

for any faithful representation G → GL(n). So, by Merkurjev’s theorem, we have

K_{0}BG = \underleftarrow{lim}_{N}Z ⊗_{R(GL(N+n))} R(GL(N) × G)
= \underleftarrow{lim}_{N}[Z ⊗_{R(GL(N+n))} R(GL(N) × GL(n))] ⊗_{R(GL(n))} R(G).

The ring in brackets is K_{0}(GL(N + n)/GL(N) × GL(n)), or equivalently K_{0} of the Grassmannian of n-planes in affine (N + n)-space. In particular, we easily compute that the representation ring of GL(n) maps onto this ring, with kernel contained in F_{γ}^{N+1}R(GL(n)). (See Atiyah [1] for the definition of the gamma filtration of a λ-ring.) It follows that the representation ring R(G) maps onto the above ring K_{0}(GL(N + n)/GL(N) × G) with kernel contained in F_{γ}^{N+1}R(G). Also, this kernel contains F_{γ}^{M}R(G) for some large M, just because GL(N +n)/GL(N)×G is finite-dimensional. It follows that K_{0}BG is isomorphic to the completion of the representation ring R(G) with respect to the gamma filtration. By Atiyah’s arguments ([1], pp. 56-57), this is the same as the completion of R(G) with respect to powers of the augmentation ideal. QED.

Esnault, Kahn, Levine, and Viehweg pointed out that Merkurjev’s theorem implies that CH^{1}BG and CH^{2}BG are generated by Chern classes of representations of G ([12], Appendix C). In general, we have the following statement.

Corollary 3.2. For any algebraic group G over a field k, the subgroup of the Chow group CH^{i}BG generated by Chern classes c_{i} of k-representations of G contains (i − 1)!CH^{i}BG.

Proof. For any smooth algebraic variety X over a field k, the algebraic Kgroup K_{0}X has two natural filtrations, the geometric filtration (by codimension of support) and the gamma filtration [17]. There are natural maps from both CH^{i}X and gr_{γ}^{i}K_{0}X to gr_{geom}^{i}K_{0}X. The map CH^{i}X → gr_{geom}^{i}K_{0}X is surjective, and the ith Chern class gives a map back such that the composition

CH^{i}X ⟶ gr_{geom}^{i}K_{0}X ⟶^{c_{i}} CH^{i}X

is multiplication by (−1)^{i−1}(i − 1)!, by Riemann-Roch without denominators [20], [17]. It follows that the surjection from CH^{i}X to gr_{geom}^{i}K^{0}X becomes an isomorphism after inverting (i−1)!, and the subgroup of the Chow group CH^{i}X generated by Chern classes c_{i} of vector bundles on X contains (i − 1)!CH^{i}X.

Apply this to the variety X = GL(N + n)/(GL(N) × G), for any faithful representation G → GL(n) and any N ≥ i. By Remark 1.4, CH^{i}BG maps isomorphically to CH^{i}X, and by Merkurjev’s theorem, K_{0}X is generated by representations of G. It follows that the subgroup of CH^{i}BG generated by Chern classes c_{i} of representations of G contains (i − 1)!CH^{i}BG. QED.

Now suppose that G is an algebraic group over the complex numbers. Then, in keeping with the observation that the homomorphism CH^{∗}BG → MU^{∗}BG⊗_{MU^{∗}} Z is often an isomorphism, we can prove properties analogous to the above property of CH^{∗}BG for MU^{∗}BG ⊗_{MU^{∗}} Z, as follows.

Theorem 3.3. For any complex algebraic group G, the subgroup of CH^{i}BG generated by Chern classes c_{i} of representations of G contains (i − 1)!CH^{i}BG, and the subgroup of (MU^{∗}BG ⊗_{MU^{∗}} Z)^{2i} generated by Chern classes c_{i} of representations contains (i − 1)!(MU^{∗}BG ⊗_{MU^{∗}} Z)^{2i}.

Moreover, there are maps

gr_{γ}^{i}R(G) ⟶^{c_{i}} CH^{i}BG ⟶ (MU^{∗}BG ⊗_{MU^{∗}} Z)^{2i} ⟶ gr_{top}^{2i}R(G)

such that the first two maps are surjective after inverting (i − 1)!. (We will give Atiyah’s definition of the topological filtration on R(G) in the proof.) The composition of all three maps is (−1)^{i−1}(i − 1)! times the natural map from gr_{γ}^{i}R(G) to gr_{top}^{2i}R(G). So, for example, if that natural map is injective, then the groups gr_{γ}^{i}R(G), CH^{i}BG, and (MU^{∗}BG ⊗_{MU^{∗}} Z)^{2i} all become isomorphic after inverting (i − 1)!.

Finally, the odd-degree group (MU^{∗}BG ⊗_{MU^{∗}} Z)^{2i+1} is killed by i!.

Proof. For any smooth algebraic variety X, the proof of Corollary 3.2 shows that the natural map

CH^{i}BG → gr_{geom}^{i}K_{0}X

and the Chern class map

c_{i} : gr_{geom}^{i}K_{0}X → CH^{i}BG

become isomorphisms after inverting (i − 1)!. In particular, both groups are generated by Chern classes of elements of K_{0}X after inverting (i−1)!. Also, by definition of the gamma filtration,the image of the natural map

gr_{γ}^{i}K_{0}X → gr_{geom}^{i}K_{0}X

is the subgroup of gr_{geom}^{i}K_{0}X generated by Chern classes c_{i} of elements of K_{0}X. It follows that this map becomes surjective, although not necessarily injective, after inverting (i − 1)!. So the composition

gr_{γ}^{i}K_{0}X ⟶ gr_{geom}^{i}K_{0}X ⟶^{c_{i}} CH^{i}X

becomes surjective after inverting (i − 1)!.

We can apply all this to the classifying space BG, viewed as a limit of smooth algebraic varieties over k, using the identification of the algebraic K-group K_{0}BG with the completed representation ring of G (Theorem 3.1). The map gr_{γ}^{i}R(G) → CH^{i}BG in the statement is defined as the composition

gr_{γ}^{i}R(G) ⟶ gr_{geom}^{i}R(G) ⟶^{c_{i}} CH^{i}BG.

It becomes surjective after inverting (i − 1)!, as we want.

We showed in Corollary 3.2 that the subgroup of CH^{i}BG generated by Chern classes c_{i} of representations of G contains (i − 1)!CH^{i}BG. Now let us prove the analogous statement for complex cobordism. We need the following “Riemann-Roch without denominators” formula for complex cobordism. The statement uses the map of cohomology theories MU^{∗}X → K^{∗}X defined by Atiyah-Hirzebruch and Conner-Floyd [2], [10]; when X is a point, the homomorphism MU^{∗} → K^{∗} is simply the Todd genus. Using the periodicity of topological K-theory, we have a map MU^{2i}X → K^{2i}X = K^{0}X.

Lemma 3.4. For any topological space X, the composition

MU^{2i}X ⟶ K^{0}X ⟶^{c_{i}} (MU^{∗}X ⊗_{MU^{∗}} Z)^{2i}

is multiplication by (−1)^{i−1}(i − 1)!.

Proof. Let us first show that this is correct after we map further from the group (MU^{∗}X ⊗_{MU^{∗}} Z)^{2i} into rational cohomology H^{2i}(X, Q). It will suffice to show that the image in K^{0}X of an element x ∈ MU^{2i}X has Chern character ch_{i} in H^{2i}(X, Q) equal to the image of x under the obvious map and has ch_{j} = 0 for j < i. This is enough because the Chern character ch_{i} of an element of K^{0}X for which ch_{j} is 0 for j < i is given by

ch_{i} = (−1)^{i−1}c_{i}/(i − 1)!.

Because complex cobordism is a connective cohomology theory (MU^{∗}(point) is 0 in positive degrees), any element x ∈ MU^{2i}X restricts to 0 on the 2j-skeleton of X for j < i, which implies that the image of x in K^{0}X has Chern character ch_{j} equal to 0 in H^{2j}(X, Q) for j < i. Also, because MU^{0}(point) is just Z, the natural transformation from MU^{2i}X to H^{2i}(X, Q) given by mapping to K^{0}X and taking the Chern character ch_{i} must be a rational multiple of the obvious map. We want to show that the rational multiple is 1. It suffices to check this when X is the pair (S^{2i}, point). Then MU^{2i}X = MU^{0} = Z via the suspension isomorphism, and since the map from complex cobordism to K-theory is multiplicative, the element 1 in this group maps to 1 in K^{2i}X when we identify this group with K^{0} = Z via the suspension isomorphism. So the calculation comes down to: if we start with the element 1 in K^{0} = Z, identify it with an element K^{2i}X (where X = (S^{2i}, point)) via the suspension isomorphism, and then identify this with K^{0}X via periodicity, show that the Chern character ch_{i} of the resulting element is 1 in H^{2i}(X, Q) = Q. This is a classical property of Bott periodicity ([3], p. 16).

To prove the above formula as stated, it suffices to prove it in the universal case, where X is the 2ith space in the Ω-spectrum MU. The point is that X has torsion-free cohomology by Wilson ([44], pp. 52-53). So MU^{∗}X ⊗_{MU^{∗}} Z is equal to H^{∗}(X, Z) (by the Atiyah-Hirzebruch spectral sequence, as in [42], p. 471) and injects into H^{∗}(X, Q). Thus the desired formula is true since it is true in rational cohomology. QED.

Applying Lemma 3.4 to the classifying space of a complex algebraic group G, we deduce that the subgroup of MU^{∗}BG ⊗_{MU^{∗}} Z in degree 2i which is generated by Chern classes c_{i} of representations of G contains (i − 1)! times the whole group. Since the topological K-group K^{1}BG is 0, applying Lemma 3.4 to the suspension of BG shows that MU^{∗}BG ⊗_{MU^{∗}} Z in odd degree 2i + 1 is killed by i!.

Finally, for any CW complex X, the map from MU^{2i}X to K^{0}X lands in F_{top}^{2i}K^{0}X, where the topological filtration of K^{0}X is defined by letting F_{top}^{2i}K^{0}X be the subgroup of K^{0}X which restricts to 0 on the (2i − 1)-skeleton of X. This is clear because every element of MU^{2i}X restricts to 0 on the (2i−1)-skeleton (because complex cobordism, unlike K-theory, is a connective cohomology theory, meaning that MU^{∗}(point) is 0 in positive degrees). The quotient group (MU^{∗}X ⊗_{MU^{∗}} Z)^{2i} maps to the associated graded group gr_{top}^{2i}K^{0}X, and it is elementary to identify the composition

gr_{γ}^{i}K^{0}X ⟶^{c_{i}} (MU^{∗}X ⊗_{MU^{∗}} Z)^{2i} ⟶ gr_{top}^{2i}K^{0}X

with (−1)^{i−1}(i − 1)! times the natural map. For X = BG, this completes the proof of Theorem 3.3, since the topological filtration on the representation ring R(G) is defined as the pullback of the topological filtration of K^{0}BG. QED.

Corollary 3.5. For any complex algebraic group G, the map from the Chow group CH^{i}BG to (MU^{∗}BG⊗_{MU^{∗}} Z)^{2i} is an isomorphism for i ≤ 2. The map from these groups to H^{2i}(BG, Z) is an isomorphism for i ≤ 1 and injective for i = 2. Also, for i ≤ 2, these groups are generated by Chern classes c_{i} of representations of G.

Proof. The map from CH^{i}BG to (MU^{∗}BG⊗_{MU^{∗}} Z)^{2i} is an isomorphism for i = 1 by Theorem 3.3, using that gr_{γ}^{1}R(G) = gr_{top}^{1}R(G) = H^{2}(BG, Z) ([1], p. 58). Unfortunately, this argument is not enough to prove the desired isomorphism for i = 2, since E. Weiss found for the group G = A_{4} that the map from gr_{γ}^{2}R(G) to gr_{top}^{2}R(G) has a kernel of order 2 (see Thomas [40]). At least Theorem 3.3 shows that the map from CH^{2}BG to (MU^{∗}BG ⊗_{MU^{∗}} Z)^{4} is surjective with finite kernel.

The point is to apply an argument of Bloch and Merkurjev-Suslin, as formulated by Colliot-Thélène ([9], p. 13). For each smooth algebraic variety X over a field k and each prime number l invertible in k, this argument gives a commutative diagram

H_{Zar}^{1}(X, H_{et}^{2}(Q_{l}/Z_{l}(i))) ⟶_{γ_{2}} H_{et}^{3}(X, Q_{l}/Z_{l}(i))
α_{2}↓  ↓
CH^{2}(X)_{l-tors}  ⟶ H_{et}^{4}(X, Z_{l}(i))

such that γ_{2} is injective and α_{2} is surjective. Take X to be a quotient variety (V −S)/G for a complex algebraic group G, with S of large codimension in V. Then H^{3}(X, Q_{l}/Z_{l}) maps injectively to H^{4}(X, Z_{l}), since H^{3}(X, Q) = H^{3}(BG, Q) = 0. It follows that CH^{2}(X)_{l-tors} injects into H^{4}(X, Z_{l}), or equivalently that CH^{2}(BG)_{l-tors} injects into H^{4}(BG, Z_{l}). Since we know that the map from CH^{2}BG to (MU^{∗}BG⊗_{MU^{∗}} Z)^{4} is surjective with finite kernel, it must be an isomorphism, and these groups must inject into H^{4}(BG, Z).

Finally, Corollary 3.2 shows that CH^{2}BG is generated by Chern classes c_{2} of representations. QED.

It follows, for example, that CH^{2}BG → H^{4}(BG, Z) = Z is not surjective, for all simply connected simple groups G over C other than SL(n) and Sp(2n). The analogous statement for G = SO(4) was the source of the examples of varieties with nonzero Griffiths group in [42].

4. Transferred Euler classes

As mentioned in the introduction, for any complex reductive group G, there is a natural class of elements of the ring MU^{∗}BG ⊗_{MU^{∗}} Z which lie in the image of the Chow ring of BG. For any i-dimensional representation E of a subgroup H ⊂ G (including G itself), E determines an algebraic vector bundle on BH (meaning an algebraic vector bundle on all the varieties (V − S)/H), so it has Chern classes in the Chow ring, in complex cobordism, and in ordinary cohomology. These are compatible under the natural homomorphisms

CH^{∗}BH → MU^{∗}BH ⊗_{MU^{∗}} Z → H^{∗}(BH, Z).

In particular, we define the Euler class χ(E) to mean the top Chern class of E in any of these groups. These homomorphisms are also compatible with transfer maps, for H of finite index in G, so the transferred Euler class tr_{H}^{G}χ(E) is an element of CH^{∗}BG which maps to the element with the same name in MU^{∗}BG ⊗_{MU^{∗}} Z and in H^{∗}(BG, Z). Sometimes we also call any Z-linear combination of transferred Euler classes a transferred Euler class. In this terminology, if G is a finite group, the transferred Euler classes form a subring of any of the above three rings, by the arguments of Hopkins, Kuhn, and Ravenel [19].

There are many finite groups G which satisfy the conjecture of Hopkins, Kuhn, and Ravenel that the Morava K-theories of BG are generated by transferred Euler classes; by Ravenel, Wilson, and Yagita, it then follows that MU^{∗}BG is generated as a topological MU^{∗}-module by transferred Euler classes [33]. Whenever G has this property, the Chow ring of BG clearly maps onto MU^{∗}BG ⊗_{MU^{∗}} Z. For such groups G, we have a computation of im(CH^{∗}BG → H^{∗}(BG, Z)) in topological terms: it is the Z-submodule of H^{∗}(BG, Z) generated by transferred Euler classes, or equivalently it is the image of MU^{∗}BG → H^{∗}(BG, Z). The conjecture has been checked for various finite groups, including abelian groups, groups of order p^{3}, the symmetric groups, and finite groups of Lie type away from the defining characteristic, by [39], [19], and [38]. The conjecture about Morava K-theory fails for the group G considered by Kriz [22]. We can still guess that the Chow ring of BG is additively generated by transferred Euler classes for all finite groups G.

As the reader has probably noticed, we could consider the transfers of arbitrary Chern classes of representations instead of just the top Chern class. But one can show that transfers of arbitrary Chern classes are Z-linear combinations of transfers of Euler classes, so that this would not give anything new. On the other hand we definitely cannot avoid mentioning transfers; that is, MU^{∗}BG and CH^{∗}BG are not generated as algebras by the Chern classes of representations of G itself, for some groups G. Chris Stretch showed in 1987 that MU^{∗}(BS_{6}) is not generated by Chern classes; we will now give his pleasant argument.

The point is that S_{6} has two subgroups which are pointwise conjugate, but not conjugate:

H_{1} = {1, (12)(34), (12)(56), (34)(56)}
H_{2} = {1, (12)(34), (13)(24), (14)(23)}

Since they are pointwise conjugate, character theory shows that any complex representation of S_{6} has a restriction to H_{1} which is isomorphic to its restriction to H_{2} (in terms of any fixed isomorphism H_{1} ≅ H_{2}). It follows that any element of MU^{∗}(BS_{6}) in the subring generated by Chern classes of representations has the same restriction to H_{1} as to H_{2}, in terms of the fixed identification H_{1} ≅ H_{2}. So it suffices to find an element of MU^{∗}(BS_{6}) which restricts differently to the two subgroups.

One such element is the transfer from S_{4} × S_{2} ⊂ S_{6} to S_{6} of the Euler class of the representation S_{4} × S_{2} → S_{4} → GL(3, C). (The last representation is the permutation representation of S_{4} minus the trivial representation.) The double coset formula gives the restriction of this transferred element to any subgroup, and one finds that the restrictions of this element of MU^{∗}BS_{6} to MU^{∗}BH_{1} and MU^{∗}BH_{2} are different, since their images in H^{∗}(BH_{1}, Z) and H^{∗}(BH_{2}, Z) are different. (This argument actually gives the stronger conclusion that the image of MU^{∗}BS_{6} → H^{∗}(BS_{6}, Z) is not generated by Chern classes.)

5. Chow groups of Godeaux-Serre varieties

For any finite group G, Godeaux and Serre constructed smooth projective varieties with fundamental group G, as quotients of complete intersections by free actions of G ([34], section 20). Atiyah and Hirzebruch showed the falsity of the Hodge conjecture for integral cohomology using some of these varieties [4]: not all the torsion cohomology classes on these varieties can be represented by algebraic cycles. So it becomes a natural problem to compute the Chow groups of these varieties. We can at least offer a conjecture. Let G be a finite group, V a representation of G over a field k, and X a smooth complete intersection defined over k in P (V) such that G acts freely on X. We call the quotient variety X/G a Godeaux-Serre variety.

Conjecture 5.1. The natural homomorphism

CH^{i}(BG × BG_{m})_{k} → CH^{i}X/G

is an isomorphism for i < dim X/2. For very general X of sufficiently high degree, this homomorphism is an isomorphism for i < dim X.

The homomorphism mentioned in the conjecture comes from the natural G × G_{m}-bundle over X/G, where the G-bundle over X/G is obvious and the G_{m}-bundle on X/G is the one which corresponds to the G-equivariant line bundle O(1) on X ⊂ P (V).

The proof that Godeaux-Serre varieties exist [34] (for every finite group G and in every dimension r) helps to show why the conjecture is plausible. Namely, let G be a finite group and V a representation of G. We can imbed the quotient variety P (V)/G in some projective space P^{N}. Let S′ be the closed subset of P (V) where G does not act freely; for any given r ≥ 0, we can choose the representation V of G such that S′ has codimension greater than r in P (V). Then we construct Godeaux-Serre varieties X/G of dimension r by intersecting P (V)/G with a general linear space in P^{N} of codimension equal to dim P (V)/G − r. The intersection will not meet S′/G and will be a smooth compact subvariety of the smooth noncompact variety (P (V) − S′)/G. We can think of X/G as a complete intersection in (P (V) − S′)/G.

Let S ⊂ V be the union of 0 with the inverse image of S′ ⊂ P (V); then (P (V) − S′)/G can also be viewed as the quotient (V − S)/(G × G_{m}). Then S has codimension in V greater than r = dim X, so

CH^{i}(BG × BG_{m}) = CH^{i}((V − S)/(G × G_{m}))

for i ≤ r. So the above conjecture would follow if we could relate the Chow groups of (V −S)/(G×G_{m}) to those of the complete intersection X/G inside it. The precise bounds in the conjecture are those suggested by the usual conjectures on complete intersections in smooth projective varieties, due to Hartshorne [18] and Nori ([30], p. 368); see also Paranjape [32], pp. 643-644. The analogous topological statement was proved in this case by Atiyah and Hirzebruch: for Godeaux-Serre varieties over the complex numbers, the natural homotopy class of maps X/G → BG × BS^{1} is r-connected, where r = dimCX ([4], p. 42).

For this conjecture to really say anything about the Chow groups of Godeaux-Serre varieties, we need to understand the Chow ring of BG × BG_{m}. We can think of BG_{m} as the infinite-dimensional projective space, which makes it easy to check that the Chow ring of BG × BG_{m} is a polynomial ring in one variable in degree 1 over the Chow ring of BG. For general finite groups G, we have no computation of the Chow ring of BG, but it tends to be computable, as shown by various results in this paper.

It remains to say what kind of evidence can be offered for the conjecture. Even for the trivial group G = 1, the conjecture is far out of reach, since in that case it amounts to the conjectures of Hartshorne and Nori for complete intersections in projective space. For example, Hartshorne’s conjecture says that the low-codimension Chow groups of a smooth complete intersection over the complex numbers are equal to Z, but for all we know, they might even be uncountable (apart from CH^{1}). Nonetheless, there are various suggestive pieces of evidence for the above conjecture. Let us work over C in what follows.

A weak sort of evidence is that the only known examples of Godeaux-Serre varieties for which CH^{∗}(X/G) → H^{∗}(X/G, Z) can be proved to be non-surjective [4] or non-injective [42] in low codimension come from groups G for which the corresponding map for BG × BG_{m} is non-surjective or non-injective.

Also, since we do not know how to prove the triviality of the low-codimension Chow groups of a complete intersection X in projective space, we might settle for trying to understand the kernel of the pullback map CH^{∗}(X/G) → CH^{∗}X. By the obvious transfer argument, this kernel is a torsion group, killed by |G|. The torsion subgroup of the Chow groups of any variety over an algebraically closed field is countable by Suslin ([37], p. 227). In that sense, the kernel of CH^{∗}(X/G) → CH^{∗}X is under better control than CH^{∗}X.

For CH^{2}, we can prove the conjecture on this kernel for varieties X/G of sufficiently large dimension:

Theorem 5.2. For any finite group G, let X be a smooth complete intersection over C on which G acts freely. For dim X sufficiently large, depending on G, the natural map

ker(CH^{2}(BG × BC^{∗}) → CH^{2}(BC^{∗})) → ker(CH^{2}(X/G) → CH^{2}X)

is an isomorphism. The first kernel is naturally identified with CH^{1}BG⊕CH^{2}BG.

Proof. By the argument of Bloch and Merkurjev-Suslin used already in section 3, the torsion subgroup of CH^{2} of a smooth projective variety Y over C maps injectively to H^{3}(Y, Q/Z) (see [9], p. 17). Taking Y to be a Godeaux-Serre variety X/G of dimension at least 4, we have H^{3}(X/G, Q) ↪ H^{3}(X, Q) = 0, and so the torsion subgroup of CH^{2}(X/G) maps injectively to the ordinary cohomology group H^{4}(X/G, Z). If X has dimension at least 5, this cohomology group is equal to H^{4}(BG′, Z), where we write G′ = G × C^{∗}. Thus the kernel of CH^{2}X/G → CH^{2}X maps injectively to H^{4}(BG′, Z). The image in cohomology of that kernel clearly contains the image of the kernel of CH^{2}BG′ → CH^{2}BC^{∗} = Z, and we will show that equality holds when X has sufficiently large dimension, depending on G.

In fact, we know by Corollary 3.5 that the map

CH^{2}BG′ → (MU^{∗}BG′ ⊗_{MU^{∗}} Z)^{4}

is an isomorphism. And if (BG)_{n} denotes the n-skeleton of BG, Lemma 2.2 (Landweber’s theorem) says that for every n there exists an r ≥ n such that

im(MU^{∗}BG′ → MU^{∗}(BG′)_{n}) = im(MU^{∗}(BG′)_{r} → MU^{∗}(BG′)_{n}).

Now, for any space X, the image of an element x ∈ MU^{4}X in H^{4}(X, Z) only depends on the restriction of x to MU^{4} of the 4-skeleton X_{4}, since H^{4}(X, Z) injects into H^{4}(X_{4}, Z). So the above statement implies that there is an r ≥ 4 such that

im(MU^{4}BG′ → H^{4}(BG′, Z)) = im(MU^{4}(BG′)_{r} → H^{4}(BG′, Z)).

Since a Godeaux-Serre variety X/G of dimension r contains the r-skeleton of BG′ up to homotopy, it follows that

im(MU^{4}BG′ → H^{4}(X/G, Z)) = im(MU^{4}X/G → H^{4}(X/G, Z))

for Godeaux-Serre varieties X/G of dimension at least r. Thus, for Godeaux-Serre varieties of sufficiently large dimension, the image of the Chow group CH^{2}X/G in integral cohomology is contained in the image of MU^{4}X/G, hence in the image of MU^{4}BG′, and hence (by Corollary 3.5) in the image of CH^{2}BG′. In particular, if we start with an element x ∈ CH^{2}(X/G) which pulls back to 0 in CH^{2}X, then, as we have said, the class of x in integral cohomology is the image of an element of CH^{2}BG′, and this element must pull back to 0 in CH^{2}BC^{∗} = H^{4}(BC^{∗}, Z) = Z because x pulls back to 0 in H^{4}(X, Z) = H^{4}(BC^{∗}, Z) = Z.

Finally, the maps of both ker(CH^{2}BG′ → Z) and ker(CH^{2}X/G → CH^{2}X) to H^{4}(X/G, Z) = H^{4}(BG′, Z) are injective. For the first group, this is Corollary 3.5. For the second group, we proved this two paragraphs back. It follows that the kernel of CH^{2}BG′ → CH^{2}BC^{∗} = Z maps isomorphically to the kernel of CH^{2}X/G → CH^{2}X. QED.

6. Some examples of the Chow Künneth formula

As a warmup to our discussion of the Chow groups of symmetric products (which we will use to compute the Chow cohomology of the symmetric group), we make some easy remarks about the Chow groups of a product in this section.

Let X and Y be algebraic varieties over a field k. We describe a special situation in which the Chow groups of X × Y are determined by the Chow groups of X and Y by the “Chow Künneth formula”:

CH_{∗}(X × Y) ≅ CH_{∗}(X) ⊗Z CH_{∗}Y.

This formula is certainly false in general, for example for the product of an elliptic curve with itself. The interesting thing is that the formula holds in some cases even where CH_{∗}X and CH_{∗}Y have a lot of torsion, whereas the corresponding formula for integer homology would have an extra “Tor” term.

Lemma 6.1. If X is any variety and Y is a variety which can be partitioned into open subsets of affine spaces, then the map

CH_{∗}X ⊗Z CH_{∗}Y → CH_{∗}(X × Y)

is surjective.

Proof. We recall the basic exact sequence for the Chow groups. Let Y be a variety, S ⊂ Y a closed subvariety, and U = Y − S. Then the sequence

CH_{∗}S → CH_{∗}Y → CH_{∗}U → 0

is exact.

We know the Chow groups of affine space: CH_{i}(A^{n}) is Z if i = n and 0 otherwise. The basic exact sequence implies that the Chow groups of a nonempty open subset Y ⊂ A^{n} are the same: CH_{∗}Y = Z in dimension n. We also know that

CH_{∗}X ⊗Z CH_{∗}A^{n} ≅ CH_{∗}(X × A^{n})

for any variety X. It follows that

CH_{∗}X ⊗Z CH_{∗}Y → CH_{∗}(X × Y)

is surjective for any open subset Y ⊂ A^{n}, since X × Y is an open subset of X × A^{n}.

Now let Y be any variety, S ⊂ Y a closed subvariety, and U = Y − S. Let X be a variety. Then we have exact sequences as shown.

CH_{∗}X ⊗Z CH_{∗}S ⟶ CH_{∗}X ⊗Z CH_{∗}Y ⟶ CH_{∗}X ⊗Z CH_{∗}U ⟶ 0
↓  ↓  ↓
CH_{∗}(X × S) ⟶ CH_{∗}(X × Y) ⟶ CH_{∗}(X × U) ⟶ 0.

It follows that if CH_{∗}X⊗CH_{∗}S → CH_{∗}(X×S) and CH_{∗}X⊗CH_{∗}U → CH_{∗}(X×U) are surjective, then so is CH_{∗}X ⊗ CH_{∗}Y → CH_{∗}(X × Y). The theorem follows. QED.

Lemma 6.2. If X and Y are varieties such that the maps CH_{∗}X → H_{∗}^{BM}(X, Z) and CH_{∗}Y → H_{∗}^{BM}(Y, Z) are split injections of abelian groups, then the map

CH_{∗}X ⊗Z CH_{∗}Y → CH_{∗}(X × Y)

is injective. If in addition X or Y can be partitioned into open subsets of affine spaces, then we have

CH_{∗}X ⊗Z CH_{∗}Y ≅ CH_{∗}(X × Y),

and X × Y also has a split injection CH_{∗}(X × Y) ↪ H_{∗}^{BM}(X × Y, Z).

Proof. Use the diagram

CH_{∗}X ⊗Z CH_{∗}Y  ⟶ CH_{∗}(X × Y)
↓  ↓
0 → H_{∗}^{BM}X ⊗Z H_{∗}^{BM}(Y, Z) ⟶ H_{∗}^{BM}(X × Y, Z),

where the bottom arrow is split injective. QED.

Fulton has asked whether the Chow Künneth formula might be true assuming only that Y can be cut into open subsets of affine spaces, with no assumption on X and no homological assumption. In particular, Fulton, MacPherson, Sottile, and Sturmfels proved this when Y is a (possibly singular) toric variety or spherical variety [15]. I generalized their result to an arbitrary “linear variety” Y in the sense of Jannsen [41].

In particular, let G_{1} be any algebraic group such that there are quotients (V − S)/G_{1} with the codimension of S arbitrarily large which are linear varieties; this is true for most of the groups G_{1} for which we can compute the Chow ring of BG_{1}. Then

CH^{∗}B(G_{1} × G_{2}) = CH^{∗}BG_{1} ⊗Z CH^{∗}BG_{2}

for all algebraic groups G_{2}. Of course, this would not be true for integral cohomology in place of the Chow ring, say for G_{1} = Z/p. But something similar is true for complex cobordism: if G_{1} is a compact Lie group such that the Morava K-theories of BG_{1} are concentrated in even degrees, then

MU^{∗}B(G_{1} × G_{2}) = MU^{∗}BG_{1} ⊗_{MU^{∗}} MU^{∗}BG_{2}

for all compact Lie groups G_{2} [33].

A special case of all this is that for all abelian groups G over C, the Chow ring of BG is the symmetric algebra on CH^{1}BG = Hom(G, C^{∗}). This ring maps isomorphically to MU^{∗}BG ⊗_{MU^{∗}} Z, by Landweber’s calculation of MU^{∗}BG for G abelian [23].

7. Chow groups of cyclic products, introduction

We construct operations from the Chow groups of a quasi-projective scheme X to the Chow groups of the pth cyclic product Z^{p}X := X^{p}/(Z/p). We need X quasiprojective in order to know that this quotient variety Z^{p}X exists; it will again be quasi-projective. From the construction it will be clear that these operations agree with the analogous operations on integral homology, as constructed by Nakaoka [29]. More precisely, Nakaoka constructed such operations which raise degree by either an even or an odd amount, and the operations we construct agree with his even operations.

Lemma 7.1. Let p be a prime number. Let X be a quasi-projective scheme over the complex numbers, and let Z^{p}X = X^{p}/(Z/p) denote the pth cyclic product of X. Then we will define operations

α_{i}^{j} : CH_{i}X → CH_{j}Z^{p}X, i + 1 ≤ j ≤ pi − 1

and

γ_{i} : CH_{i}X → CH_{pi}Z^{p}X, i > 0,

with the following properties. (The formulas use the natural map (CH_{∗}X)^{⊗p} → CH_{∗}X^{p} → CH_{∗}Z^{p}X.)

(1) α_{i}^{j} is a homomorphism of abelian groups.

(2) pα_{i}^{j} = 0.

(3)

γ_{i}(x + y) = γ_{i}x + ∑ α_{1} ⊗ ⋯ ⊗ α_{p} + γ_{i}y,

where the sum is over a set of representatives α = (α_{1}, ⋯, α_{p}) for the action of Z/p on the set {x, y}^{p} − {(x, ⋯, x), (y, ⋯, y)}.

(4) pγ_{i}x = x^{⊗p}.

In the statement we assumed that X was a scheme over the complex numbers. The proof will in fact construct operations with all the same properties over any field k of characteristic ≠ p which contains the pth roots of unity; the only complication is that the operations α_{i}^{j} are naturally viewed as maps:

α_{i}^{j} : CH_{i}X ⊗ μ_{p}(k)^{⊗(pi−j)} → CH_{j}Z^{p}X.

Proof. Let C be any algebraic cycle of dimension i > 0 on X. (By definition, an algebraic cycle on X is simply a Z-linear combination of irreducible subvarieties of X.) There is an associated cycle C^{p} on the product X^{p}. Since the subvarieties of X occurring in C have dimension i > 0, it is easy to see that none of the subvarieties occurring in the cycle C^{p} is contained in the diagonal Δ_{X} ≅ X ⊂ X^{p}. Since the cycle C^{p} is invariant under the permutation action of the group Z/p on X^{p}, the restriction of the cycle C^{p} to X^{p} − X (where Z/p acts freely) is the pullback of a unique cycle on Z^{p}X − X under the etale morphism X^{p} − X → Z^{p}X − X. Define the cycle Z^{p}C on Z^{p}X to be the closure of this cycle on Z^{p}X − X. Define the operation γ_{i} on a cycle C on X to be the class of this cycle Z^{p}C on Z^{p}X. We will show below that this operation is well-defined on rational equivalence classes.

First we define the other operations α_{i}^{j}. Let C be an algebraic cycle on X of dimension i > 0. Let D denote the support of C, that is, the union of the subvarieties of X that occur in the cycle C. There is an obvious Z/p-principal bundle over Z^{p}D − D. Given a pth root of unity and thus a homomorphism Z/p → k^{∗}, we get a line bundle L over Z^{p}D − Δ_{D}. (For k = C, we use the pth root of unity e^{2πi/p}.) Clearly pc_{1}L = 0 ∈ Pic(Z^{p}D − Δ_{D}). Line bundles act on the Chow groups, lowering degree by 1, so we can consider the element

c_{1}(L)^{pi−j}[Z^{p}C]

of CH_{j}(Z^{p}D − Δ_{D}) for i + 1 ≤ j ≤ pi − 1. Here Z^{p}C is the cycle defined in the previous paragraph; this cycle is supported in the closed subset Z^{p}D ⊂ Z^{p}X, and the formula refers to its restriction to a cycle on Z^{p}D − D. The elements defined by this formula are killed by p since c_{1}(L) is. Since these classes are above the dimension of Δ_{D}, they uniquely determine elements

Y_{j} ∈ CH_{j}Z^{p}D,

i + 1 ≤ j ≤ pi − 1. Clearly pY_{j} = 0 in CH_{j}Z^{p}D since this is true after restricting to the open subset Z^{p}D − D. We define the operation α_{i}^{j} on the cycle C to be the image of this class Y_{j} in CH_{i}Z^{p}X.

Let us show that the elements α_{i}^{j}C and γ_{i}C in the Chow groups of Z^{p}X are well-defined when we replace the cycle C by a rationally equivalent cycle. We have to show that for every (i + 1)-dimensional irreducible subvariety W ⊂ X × P^{1} with the second projection not constant, if we let C_{0} and C_{∞} denote the fibers of W over 0 and ∞ in P^{1}, viewed as cycles on X, then α_{i}^{j}(C_{0}) = α_{i}^{j}(C_{∞}) and γ_{i}(C_{0}) = γ_{i}(C_{∞}) in the Chow groups of Z^{p}X.

To see this, we use the fiber product W^{p}/P^{1}. The group Z/p acts on this fiber product in a natural way, acting as the identity on P^{1}, and we call the quotient scheme Z^{p}W/P^{1}. The fibers over 0 and infinity, as cycles on Z^{p}X, are exactly the cycles Z^{p}C_{0} = γ_{i}(C_{0}) and Z^{p}C_{∞} = γ_{i}(C_{∞}). Thus we have a rational equivalence between the cycles γ_{i}(C_{0}) and γ_{i}(C_{∞}), so that the operation γ_{i} is well-defined on Chow groups.

We proceed to show that the other operations α_{i}^{j} are also well-defined on Chow groups. Let j be any integer with i + 1 ≤ j ≤ pi − 1. As in the definition of α_{i}^{j}, fix a pth root of unity in k and thus a homomorphism Z/p → k^{∗}. This gives a line bundle L on Z^{p}W/P^{1}−W. Using the action of line bundles on Chow groups, we get a j + 1-dimensional Chow class on Z^{p}W/P^{1} − W associated to the action of c_{1}(L)^{r} on the fundamental cycle of the scheme Z^{p}W/P^{1} − W, where we let r = pi − j. Let Y_{j} be the closure in Z^{p}W/P^{1} of any cycle on Z^{p}W/P^{1} − W representing this class. The fibers of the cycle Y_{j} over 0 and ∞ are j-dimensional cycles on Z^{p}W_{0} and Z^{p}W_{∞}, respectively, whose restrictions to Z^{p}W_{0} − W_{0} and Z^{p}W_{∞} − W_{∞} are rationally equivalent (on these open subsets) to the cycles α_{i}^{j}(C_{0}) and α_{i}^{j}(C_{∞}). But W_{0} and W_{∞} have dimension i, which is less than the dimension j of these cycles. So in fact the fibers of Y_{j} over 0 and ∞ are rationally equivalent to the cycles α_{i}^{j}(C_{0}) and α_{i}^{j}(C_{∞}) on the whole sets Z^{p}W_{0} and Z^{p}W_{∞}. By mapping Y_{j} into X, it follows that α_{i}^{j}(C_{0}) and α_{i}^{j}(C_{∞}) are rationally equivalent on X, as we want.

We now prove properties (1)-(4) of the operations α_{i}^{j} and γ_{i}. It is easy to prove properties (1) and (3), describing what happens to the operations after multiplying by p. Namely γ_{i}[C] is represented by the cycle Z^{p}C on Z^{p}X, and p times this cycle is the image of the cycle C^{p} on X^{p}, since the map C^{p} → Z^{p}C is generically p to

1. (We are using that i = dim C is greater than 0 here.) And the classes α_{i}^{j}C are obviously p-torsion, since the Chern class c_{1}L used to define them is p-torsion.

It is not much harder to describe what γ_{i} of a sum of two cycles is. We have

Z^{p}(C_{1} + C_{2}) = Z^{p}C_{1} + ∑ f_{∗}(α_{1} × ⋯ × α_{p}) + Z^{p}C_{2},

where f denotes the map X^{p} → Z^{p}X and the sum runs over some set of representatives α for the orbits of the group Z/p on the set {C_{1}, C_{2}}^{p}−{(C_{1}, ⋯, C_{1}), (C_{2}, ⋯, C_{2})}. This immediately implies the formula for γ_{i}(x + y).

Finally, the formula above implies an equality in CH_{∗}(Z^{p}X − X):

c_{1}(L)^{r}[Z^{p}(C_{1} + C_{2})] = c_{1}(L)^{r}[Z^{p}C_{1}] + ∑ c_{1}(L)^{r}f_{∗}[α_{1} × ⋯ × α_{p}] + c_{1}(L)^{r}[Z^{p}C_{2}],

for all r ≥ 0. For r ≥ 1, since f^{∗}c_{1}(L) = 0, the terms in the middle are all 0, so we have that

α_{i}^{j}(x + y) = α_{i}^{j}x + α_{i}^{j}y.

To be honest, we have only checked this equality in CH_{∗}(Z^{p}X − X). But we can apply this argument with X replaced by its subscheme C_{1} ∪ C_{2}. Then we have proved that

α_{i}^{j}(x + y) = α_{i}^{j}x + α_{i}^{j}y

in the Chow groups of Z^{p}(C_{1} ∪ C_{2}) − (C_{1} ∪ C_{2}). Since these elements of the Chow groups have dimension greater than that of C_{1} ∪ C_{2}, we have the same equality in the Chow groups of Z^{p}(C_{1} ∪ C_{2}). This implies the equality in the Chow groups of Z^{p}X. QED.

We can sum up what we have done by defining a certain functor F_{p} from graded abelian groups A_{∗} to graded abelian groups; Lemma 7.1 amounts to the assertion that for any variety X there is a natural map F_{p}CH_{∗}X → CH_{∗}Z^{p}X. Namely, let F_{p}A_{∗} be the graded abelian group generated by A_{∗} ⊗Z ⋯ ⊗Z A_{∗} (p copies of A_{∗}) together with μ_{p}(k)^{⊗(pi−j)} ⊗Z A_{i} in degree j for i + 1 ≤ j ≤ pi − 1, and elements γ_{i}x_{i} in degree pi, where x ∈ A_{i} and i > 0. To get F_{p}A_{∗} we divide out by the relations

x_{1} ⊗ ⋯ ⊗ x_{p} = x_{2} ⊗ ⋯ ⊗ x_{p} ⊗ x_{1}
pγ_{i}x = x^{⊗p}
γ_{i}(x + y) = γ_{i}x + ∑ α_{1} ⊗ ⋯ ⊗ α_{p} + γ_{i}y

Here, as before, the sum in the last formula runs over the Z/p-orbits α in the set {x, y}^{p} − {x, y}.

8. Chow groups of cyclic products, concluded

We can now calculate the Chow groups of cyclic products for the same class of varieties for which we proved the Chow Künneth formula (about ordinary products). Again we work over C, although the same proofs work for varieties over any field k of characteristic ≠ p which contains the pth roots of unity, using etale homology in place of ordinary homology. Actually, the following Lemma is stated in terms of Borel-Moore homology, which can be defined as the homology of the chain complex of locally finite singular chains on a locally compact topological space [13].

Since we cannot even compute the Chow groups of a product in general, as discussed in section 6, we cannot expect to compute the Chow groups of arbitrary symmetric products. Fortunately we only need to analyze symmetric products of a special class of varieties in order to compute the Chow ring of the symmetric group.

Lemma 8.1. (1) There is a functor F_{p}, defined in section 7, from graded abelian groups to graded abelian groups, such that there is a natural map F_{p}CH_{∗}X → CH_{∗}Z^{p}X for any quasi-projective variety X.

(2) If X can be cut into open subsets of affine spaces, then so can Z^{p}X, and the map F_{p}CH_{∗}X → CH_{∗}Z^{p}X is surjective.

(3) If CH_{∗}X → H_{∗}^{BM}(X, Z) is split injective, then the composition F_{p}CH_{∗}X → CH_{∗}Z^{p}X → H_{∗}^{BM}(X, Z) is split injective.

Thus if X satisfies hypotheses (2) and (3), then so does Z^{p}X, and F_{p}CH_{∗}X ≅ CH_{∗}Z^{p}X.

Here (1) was proved at the end of the last section.

Proof. (2) This follows from the observations that the cyclic product Z^{p}A^{i} can be cut into a union of open subsets of affine spaces of dimensions i + 1, i + 2, …, pi, corresponding to the operations α_{i}^{j} and γ_{i} on the fundamental class of A^{n}, and that an open subset of affine space has Chow groups equal to Z in the top dimension and 0 below that.

To prove this description of Z^{p}A^{i}, we can analyze, more generally, any quotient variety V/(Z/p), where V is a representation of Z/p over a field k in which p is invertible and which contains the pth roots of unity. Suppose V is a nontrivial representation of Z/p. Then we can write V = W ⊕ L where L is a nontrivial 1-dimensional representation of Z/p. The quotient variety V/(Z/p) can be cut into W/(Z/p) and a W -bundle over L^{⊗p} − 0 ≅ A^{1} − 0. The latter bundle can be seen, for example by direct calculation, to be isomorphic to W × (A^{1} − 0) as a variety. This analysis gives by induction a decomposition of V/(Z/p) into open subsets of affine spaces, as we need. QED.

Proof. (3) Suppose that CH_{∗}X → H_{∗}^{BM}(X, Z) is split injective. Write the finitely generated abelian group CH_{∗}X as a direct sum of cyclic groups (Z/a_{i}) · e^{i}, where a_{i} is 0 or a prime power, e^{i} ∈ CH_{∗}X, and 1 ≤ i ≤ n. Let S ⊂ {1, …, n} be the set of i such that a_{i} is 0 or a power of p and dim e^{i} > 0. Then, from the definition of the functor F_{p}, it is easy to check that F_{p}CH_{∗}X is a quotient of the group

A_{∗} := ⊕Z/(a_{i_{1}}, …, a_{i_{p}}) · e^{i_{1}} ⊗ ⋯ ⊗ e^{i_{p}} ⊕ ⊕_{i}Z/(pa_{i}) · γ(e^{i}) ⊕ ⊕_{i,j}Z/p · α^{j}(e^{i}).

Here the first sum runs over a set of orbit representatives for the action of Z/p on {1, …, n}^{p} − S, the second sum is over i ∈ S, and the second sum is over i ∈ S and dim e^{i} + 1 ≤ j ≤ p dim e^{i} − 1.

We will show that the map A_{∗} → H_{∗}^{BM}(Z^{p}X, Z) is split injective. This will imply that F_{p}CH_{∗}X is actually isomorphic to A_{∗}, not just a quotient of it, and also that the composition F_{p}CH_{∗}X → CH_{∗}Z^{p}X → H_{∗}^{BM}(Z^{p}X, Z) is split injective, thus proving (3).

We use the compatibility (which is clear) of our operations γ_{i} and α_{i}^{j} on Chow groups with the even-degree operations from the homology of X to the homology of Z^{p}X defined by Nakaoka [29]. From Nakaoka’s basis for the homology of Z^{p}X in terms of his operations, we see that F_{p}(H_{∗}^{BM}(X, Z)) maps by a split injection into H_{∗}^{BM}(Z^{p}X, Z). Since CH_{∗}X is a direct summand of H_{∗}^{BM}(X, Z), we deduce that A_{∗} = F_{p}CH_{∗}X maps by a split injection into H_{∗}^{BM}(Z^{p}X, Z). QED.

9. Wreath products

Lemma 9.1. Let G be a finite group which is an iterated wreath product Z/p ≀ Z/p ≀ ⋯ ≀ Z/p. Then the natural homomorphism from the Chow ring of (BG)C to the integer cohomology ring of BG is injective, and in fact additively split. The same is true for products of groups of this form.

Proof. It follows from Lemma 8.1 that if G is a finite group such that BG can be approximated by smooth quasi-projective varieties satisfying conditions (2) and (3) in the lemma, then the wreath product Z/p ≀ G has the same property. That is, writing out conditions (2) and (3) in more detail, the wreath product Z/p≀G has the property that B(Z/p≀G) can be approximated by smooth varieties which can be cut into open subsets of affine spaces, and the map CH^{∗}B(Z/p≀G) → H^{∗}(B(Z/p≀G), Z) is split injective. In deducing this statement from Lemma 8.1, the point is that if X is a smooth variety which approximates BG up to some high dimension, then the complement of the diagonals in the cyclic product Z^{p}X is a smooth variety which approximates B(Z/p ≀ G), and those diagonals have high codimension, so that the Chow cohomology in low degrees of Z^{p}X − (diagonals) is equal to the low-codimension Chow groups of the singular variety Z^{p}X; and those groups are what Lemma 8.1 computes.

This implies the lemma for wreath products of copies of Z/p, by induction. For products of such groups, the lemma follows, using the Chow Künneth formula (Lemma 6.2). QED.

Lemma 8.1 actually gives an explicit calculation of the Chow cohomology groups of BG, for groups G as in Lemma 9.1. In particular, we can see from this calculation that for such groups G, the mod p Chow ring of the wreath product Z/p ≀ G is detected on the two subgroups G^{p} and Z/p × G. This means that the homomorphism

CH^{∗}B(Z/p ≀ G)/p → CH^{∗}(BG^{p})/p ⊕ CH^{∗}B(Z/p × G)/p

is injective. Applying this repeatedly, we find that the mod p Chow ring of a group G as in Lemma 9.1 is detected on the elementary abelian subgroups of G.

10. The symmetric group: injectivity

The calculation of CH^{∗}BS_{n}, locally at the prime p (that is, the calculation of CH^{∗}BS_{n} ⊗Z Z_{(p)}), follows from the calculation of CH^{∗}B(Z/p ≀ ⋯ ≀ Z/p). Namely, let H be the p-Sylow subgroup of the symmetric group S_{n}; if one writes n = p^{i_{1}} + p^{i_{2}} + ⋯, i_{1} < i_{2} < ⋯, then H = (Z/p)^{≀i_{1}} × (Z/p)^{≀i_{2}} × ⋯, where (Z/p)^{≀i} denotes the i-fold wreath product Z/p ≀ ⋯ ≀ Z/p. We know the Chow cohomology groups of BH by Lemma 9.1 and the comments afterward. In particular we know that the Chow cohomology groups of BH inject onto a direct summand of its integer cohomology. The p-localization of the Chow groups of BS_{n}, i.e., CH^{∗}(BS_{n})_{(p)} := CH^{∗}(BS_{n}) ⊗Z Z_{(p)}, can then be described as follows.

Lemma 10.1.

CH^{∗}(BS_{n})_{(p)} = CH^{∗}BH ∩ H^{∗}(BS_{n}, Z)_{(p)} ⊂ H^{∗}(BH, Z)_{(p)}.

In particular, the Chow ring of the symmetric group maps injectively to its integral cohomology, by applying this lemma for each prime number p. In fact the proof will show that this map is a split injection.

Proof. A standard transfer argument shows that if H denotes the p-Sylow subgroup of a finite group G, then we have a split injection

H^{∗}(BG, Z)_{(p)} ↪ H^{∗}(BH, Z),

and the image is precisely the “G-invariant” subgroup of H^{∗}BH ([7], p. 84). Here we say that x ∈ H^{∗}BH is G-invariant if x|_{H∩gHg^{−1}} = gxg^{−1}|_{H∩gHg^{−1}} for all g ∈ G. The same argument gives an injection of Chow rings:

CH^{∗}(BG)_{(p)} ↪ CH^{∗}BH

is the subgroup of G-invariant elements of CH^{∗}BH.

Thus the inclusion CH^{∗}(BG)_{(p)} ⊂ CH^{∗}BH ∩ H^{∗}BG_{(p)} is obvious. To prove the opposite direction, we have to show that an element of CH^{∗}BH whose image in H^{∗}BH is G-invariant is itself G-invariant. This is not obvious, since we do not know whether CH^{∗}B(H ∩ gHg^{−1}) injects into H^{∗}B(H ∩ gHg^{−1}) even though we know this for H itself. But in fact it is easy. Let x ∈ CH^{∗}BH have G-invariant image in H^{∗}BH. Then the standard transfer argument, already used above, shows that res_{H}^{G}tr_{H}^{G}x = (G : H)x ∈ H^{∗}BH. But the restriction maps in CH^{∗} and H^{∗} are compatible, and likewise for transfer. And CH^{∗}BH injects into H^{∗}BH, so we actually have res_{H}^{G}tr_{H}^{G}x = (G : H)x ∈ CH^{∗}BH. Since (G : H) is a p-local unit, this implies that x is the restriction of an element of CH^{∗}BG and so is G-invariant. QED.

The same transfer argument shows that CH^{∗}(BS_{n})⊗Z/p injects into H^{∗}(BS_{n}, Z/p). It follows that the map CH^{∗}(BS_{n})_{(p)} → H^{∗}(BS_{n}, Z)_{(p)} is a split injection.

11. The Chow ring of the symmetric group

Theorem 11.1. The homomorphism

CH^{∗}BS_{n} → MU^{∗}BS_{n} ⊗_{MU^{∗}} Z

is an isomorphism. These rings map by an additively split injection to the cohomology ring H^{∗}(BS_{n}, Z).

Proof. In section 9, we showed that the Chow ring of BS_{n} maps additively split injectively to its cohomology. A fortiori, the Chow ring injects into MU^{∗}BS_{n}⊗_{MU^{∗}} Z. This injection is in fact an isomorphism, since Hopkins-Kuhn-Ravenel [19] and Ravenel-Wilson-Yagita [33] have shown that the topological MU^{∗}-module MU^{∗}BS_{n} is generated by transferred Euler classes, and the image of such a class in the quotient ring MU^{∗}BS_{n} ⊗_{MU^{∗}} Z comes from CH^{∗}BS_{n}. QED.

This proof shows that the Chow ring of any symmetric group is generated by transferred Euler classes. That can also be seen more directly by making our computation of the Chow ring of wreath products more explicit. Indeed, the Chow ring of iterated wreath products of Z/p is generated by transferred Euler classes, and this implies the same statement for the symmetric group.

12. The ring CH^{∗}(BS_{n}) ⊗ Z/2

We now describe the ring CH^{∗}(BS_{n}) ⊗ Z/2 more explicitly.

Proposition 12.1. For the symmetric group S_{n}, the two maps

Frob : H^{i}(BS_{n}, Z/2) → H^{2i}(BS_{n}, Z/2)

(x ↦ x^{2}) and

CH^{i}(BS_{n}) ⊗ Z/2 → H^{2i}(BS_{n}, Z/2)

are injective ring homomorphisms with the same image. This implies that we have a ring isomorphism between H^{∗}(BS_{n}, Z/2) and CH^{∗}(BS_{n}) ⊗ Z/2, with H^{i} corresponding to CH^{i}.

Proof. The first homomorphism is injective because the ring H^{∗}(BS_{n}, Z/2) has no nilpotents; this in turn follows from the Z/2-cohomology of these groups being detected by elementary abelian subgroups ([25], pp. 53-57). The second homomorphism is injective by Theorem 11.1.

Next, by the calculation of the Z/2-cohomology of wreath products of copies of Z/2, H^{∗}(BS_{n}, Z/2) is generated by transferred Euler-Stiefel-Whitney classes, that is, transfers of top Stiefel-Whitney classes of real representations of subgroups of S_{n}. Then observe that the squaring map in Z/2-cohomology,

Frob : H^{i}(X, Z/2) → H^{2i}(X, Z/2),

coincides for each i with a certain Steenrod operation, thus a stable operation, which therefore commutes with transfers for finite coverings. That is, for a finite covering f : X → Y and x ∈ H^{∗}(X, Z/2), f_{∗}(x^{2}) = (f_{∗}x)^{2}. It follows that H^{∗}(BS_{n}, Z/2)^{2} is generated by transferred Euler classes of complex representations of subgroups, since for a real representation E of a group H, χ(E ⊗ C) = χ(E)^{2} ∈ H^{∗}(BH, Z/2). Thus we have the inclusion

H^{∗}(BS_{n}, Z/2)^{2} ⊂ im(CH^{∗}(BS_{n}) ⊗ Z/2 → H^{∗}(BS_{n}, Z/2)).

To get the inclusion the other way, it suffices to check that CH^{∗}(BS_{n}) ⊗ Z/2 is generated by transferred Euler classes of complexified real representations. We showed in section 11 that for any prime p, CH^{∗}(BS_{n}) ⊗ Z/p is generated by transferred Euler classes of complex representations. These representations are obtained by an inductive procedure starting from the basic representation Z/p → C^{∗}. But for p = 2, this representation is the complexification of a real representation, and so the same applies to all the resulting representations. QED.

This strange isomorphism has the property that the Stiefel-Whitney classes of the standard representation S_{n} → GL(n, R) correspond to the Chern classes of the standard representation S_{n} → GL(n, C). Just as the ring CH^{∗}BS_{6} is not generated by Chern classes, the ring H^{∗}(BS_{6}, Z/2) is not generated by Stiefel-Whitney classes.

13. The Chow ring of the symmetric group over a general field

Let G be a finite group. We can view BG as a limit of varieties defined over Q, or even over Z. By base change we get (BG)_{k} for any field k. So far we have computed CH^{∗}(BG)_{k,(p)} when G is the symmetric group and k is a field of characteristic ≠ p which contains the pth roots of unity. In this section we show that the Chow ring ⊗Z_{(p)} of the symmetric group is the same for all fields k of characteristic ≠ p.

This result can be divided into two parts. First, for any finite group G there is an action of Gal(\overline{k}/k) on CH^{∗}(BG)\overline{k}, and for some groups such as the group of order p and the symmetric group, we find that the natural map

CH^{∗}(BG)_{k} → CH^{∗}(BG)Gal(\overline{k}/k) ; \overline{k}

is an isomorphism. Then for the symmetric group, but not for the group of order p when p ≥ 3, we show that Gal(\overline{k}/k) acts trivially on CH^{∗}BG.

Even the first statement is not true for arbitrary finite groups. Even for Ktheory, which is simpler than the Chow groups, and even if the characteristic of k is prime to the order of G, the map

K_{0}BG_{k} → (K_{0}BG\overline{k})^{Gal(\overline{k}/k)}

is not always an isomorphism, although it is an isomorphism modulo torsion. The map fails to be an isomorphism if and only if G has a representation over k with Schur index not equal to 1 [35], for example when G is the quaternion group and k = R.

When G is the symmetric group, all Schur indices are equal to 1, and in fact all representations of G are defined over Q. (We will not use these facts in what follows.) In particular, the action of Gal(\overline{Q}/Q) on K_{0}BS_{n} is trivial. It follows that the action of Gal(\overline{Q}/Q) on CH^{∗}BS_{n} fixes all Chern classes of representations of S_{n}; but we need a different argument, below, to show that the Galois action on all of CH^{∗}BS_{n} is trivial.

For any finite group G, we can analyze the Galois action on CH^{∗}(BG)_{(p)} using the cycle map

CH^{i}(BG)_{\overline{k},(p)} → H^{2i}(BG\overline{k}, Z_{p}(i)),

which is Galois-equivariant. For the groups G we consider, this homomorphism is injective, so it suffices to describe the Galois action on the group H^{2i}(BG\overline{k}, Z_{p}(i)) = H^{2i}(BG\overline{k}, Z_{p}) ⊗_{Z_{p}} Z_{p}(i). The answer is simple: for all finite groups G, Gal(\overline{k}/k) acts trivially on H^{∗}(BG, Z_{p}). This is more or less obvious from the definition of etale cohomology, since we can view BG as a limit of quotient varieties X/G in which X is defined over k and every element of G acts by a map of varieties over k.

Thus the Galois action on H^{2i}(BG\overline{k}, Z_{p}(i)) just comes from the Galois action on Z_{p}(i). In the extreme case where k = Q, we can describe the Galois-fixed subgroup of H^{2i}(BG, Z_{p}(i)) as follows, by a simple calculation given by Grothendieck [16]:

H^{2i}(BG, Z_{p}(i))^{Gal(\overline{Q}/Q)} = \begin{cases}0 & if i ≢ 0 (mod p − 1) \\ ker p & if i = a(p − 1), (a, p) = 1 \\ ker p^{r+1} & if i = ap^{r}(p − 1), (a, p) = 1 \end{cases}

The subgroup fixed by Gal(\overline{Q}/Q(μ_{p})) is given by the same restrictions except that i is not required to be a multiple of p − 1.

Example 13.1. The cyclic group of order p gives an interesting example of the Galois action on CH^{∗}BG. Namely, let \overline{Q} be the algebraic closure of Q. An element of CH^{1}(BZ/p)\overline{Q} is given by a homomorphism Z/p → \overline{Q}^{∗}, so that

CH^{1}(BZ/p)\overline{Q} ≅ μ_{p}(\overline{Q}), and this description is natural with respect to automor-

phisms of \overline{Q}. Since the ring CH^{∗}(BZ/p)\overline{Q} is generated by CH^{1}, we have a nat-

ural isomorphism CH^{i}(BZ/p)\overline{Q} ≅ (μ_{p}(\overline{Q}))^{⊗i}. It follows that if we view BZ/p as a limit of varieties over Q, say, then the Chow ring of (BZ/p)Q maps to 0 except in dimensions ≡ 0 (mod p − 1). The point is that there is a natural isomorphism (μ_{p})^{⊗(p−1)} ≅ Z/p, because there is a natural generator for the cyclic group (μ_{p})^{⊗(p−1)}: take the (p − 1)st power of any generator of the cyclic group μ_{p}. This argument gives a canonical generator for (μ_{p})^{⊗i} whenever i ≡ 0 (mod p − 1), but otherwise there is none.

As a matter of fact the Chow ring of (BZ/p)Q is exactly the subring of elements of dimension ≡ 0 (mod p − 1), that is, a Z/p-polynomial algebra on one generator of degree p − 1. As it happens this is also the image of the Chow ring of BS_{p} in the Chow ring of its p-Sylow subgroup, Z/p.

We know that the Chow ring ⊗Z_{(p)} of the symmetric group over Q(μ_{p}) is the same as over \overline{Q}. But the extension Q(μ_{p})/Q has degree p − 1 which is prime to p. So the usual transfer argument for the etale covering

\begin{matrix}(BS_{n})_{Q(μ_{p})} \\ (Z/p)^{∗} & ↓ \\ (BS_{n})Q \end{matrix}

shows that CH^{∗}(BS_{n})_{Q,(p)} ≅ CH^{∗}(BS_{n})_{Q(μ_{p}),(p)}^{(Z/p)^{∗}}. So we just have to show that

the action of Gal(\overline{Q}/Q) through its quotient (Z/p)^{∗} on CH^{∗}(BS_{n})_{Q(μ_{p}),(p)} is trivial.

But our earlier comments on the Galois action on H^{2i}(BG, Z_{p}(i)), together with the injectivity of CH^{i}BS_{n} → H^{2i}(BG, Z_{p}(i)), shows that (Z/p)^{∗} acts trivially on CH^{∗}(BS_{n})_{(p)} if these p-local Chow groups are concentrated in dimensions ≡ 0 (mod p − 1).

To prove this, we use that CH^{∗}(BS_{n})/p is detected on certain explicitly known elementary abelian subgroups of S_{n}. Indeed, this mod p Chow ring injects into the mod p Chow ring of the p-Sylow subgroup of S_{n}, which is a product of iterated wreath products of copies of Z/p, and we proved in section 9 that, for a class of groups G including those considered here, the mod p Chow ring of the wreath product of Z/p ≀ G is detected on the two subgroups G^{p} and Z/p × G.

Then we observe that for all of these elementary abelian subgroups H = (Z/p)^{k} of S_{n}, the normalizer in S_{n} contains a group (Z/p)^{∗} which acts in the obvious way (by scalar multiplication) on H. It follows that (Z/p)^{∗} acts in the natural way on the polynomial ring CH^{∗}BH = Z/p[x_{1}, …, x_{k}], so that the invariants of (Z/p)^{∗} in this polynomial ring is precisely the subring consisting of elements of dimension a multiple of p − 1. But the homomorphism of the Chow ring of S_{n} into that of H automatically maps into this ring of invariants. Combining this with the previous paragraph, we have proved that the mod p Chow ring of S_{n} is nonzero only in dimensions a multiple of p − 1. It follows that the p-local Chow ring of S_{n} is also concentrated in these dimensions.

This is what we needed, three paragraphs ago, to complete the proof that the Chow groups of the symmetric group are the same over all fields of characteristic 0.

We only need a few more words to explain why the p-local Chow groups of the symmetric group are the same over any field of characteristic ≠ p. Suppose l is a prime number not equal to p. We have shown that the Chow groups of BS_{n} are the same over any field which contains F_{l}(μ_{p}), and in fact the same as in characteristic 0. The Galois group Gal(F_{l}(μ_{p})/F_{l}) is a subgroup of (Z/p)^{∗}, and since all of (Z/p)^{∗} acts trivially on the Chow groups of S_{n}, so does this subgroup. (All we are using here is that the p-local Chow groups of (BS_{n})_{F_{l}(μ_{p})} are concentrated in dimensions ≡ 0 (mod p − 1).) We deduce that the Chow groups CH^{∗}(BS_{n})_{(p)} are the same over all fields of characteristic ≠ p.

14. Generators for the Chow ring of BG

In this section, we give a simple upper bound for the degrees of a set of generators for the Chow ring of BG, for any algebraic group G (not necessarily connected). As mentioned in the introduction, nothing similar is known for the ordinary cohomology or the complex cobordism of BG.

Theorem 14.1. Let G be an algebraic group over a field, and let G ↪ H be an imbedding of G into a group H which is a product of the groups GL(n) for some integers n. Then the Chow ring CH^{∗}BG is generated as a module over CH^{∗}BH by elements of degree at most dim H/G. Here CH^{∗}BH is just a polynomial ring over Z generated by the Chern classes (see section 15).

This follows from the more precise statement:

Proposition 14.2. In the situation of the theorem, we have

CH^{∗}(H/G) = CH^{∗}BG ⊗_{CH^{∗}BH} Z.

Indeed, Proposition 14.2 implies that the Chow ring of BG maps onto that of H/G, and that CH^{∗}BG is generated as a CH^{∗}BH-module by any set of elements of CH^{∗}BG which restrict to generators for the Chow ring CH^{∗}(H/G). So, in particular, CH^{∗}BG is generated as a CH^{∗}BH-module by elements of degree at most dim H/G, thus proving Theorem 14.1. Proposition 14.2 says, more generally, that to find generators for the Chow ring of BG, it suffices to find generators for the Chow ring of a single quotient variety GL(n)/G; this will be applied in the next section to compute the Chow ring of certain classifying spaces.

Proof. (Proposition 14.2) The point is that there is a fibration

\xymatrix{H/G \ar[r] & BG \ar[d] \\  & BH}

with structure group H. (Start with the universal fibration H → EH → BH, and then form the quotient (EH ×H/G)/H; this fibers over BH with fiber H/G, and it can be identified with BG.) To avoid talking about anything infinite-dimensional, it suffices to consider the corresponding fibration

\xymatrix{H/G \ar[r] & (V − S)/G \ar[d] \\  & (V − S)/H}

associated to a representation V of H. Later in the proof we will restrict ourselves to a special class of representations of H, but since these include representations with S of arbitrarily large codimension, that will be enough to prove the proposition.

Since H is on Grothendieck’s list of “special” groups, every principal H-bundle in algebraic geometry is Zariski-locally trivial [8]. So the above fibration, being associated to the principal H-bundle over (V − S)/H, is Zariski-locally trivial.

This implies that the restriction map CH^{∗}(V −S)/G → CH^{∗}(H/G) associated to the fiber over a point x ∈ (V −S)/H is surjective. Indeed, choose a trivialization of the above bundle over a Zariski open neighborhood U of x in (V − S)/H. Then, for any subvariety Z of the fiber over x, we can spread it out to a variety Z × U, and take the closure to get a subvariety of (V − S)/G. This gives a subvariety of (V − S)/G which restricts to the given subvariety Z of the fiber H/G, so that CH^{∗}(V − S)/G → CH^{∗}(H/G) is surjective. Clearly elements of CH^{∗}(V − S)/H of positive degree restrict to 0 in the fiber H/G, so we have a surjection

CH^{∗}(V − S)/G ⊗_{CH^{∗}(V −S)/H} Z → CH^{∗}(H/G).

The proposition follows if we can prove that this map is an isomorphism. To see this, suppose that x ∈ CH^{∗}(V −S)/G restricts to 0 in the chosen fiber H/G; we have to show that x can be written as a finite sum x = ∑ a_{i}x_{i} with a_{i} ∈ CH^{>0}(V −S)/H, x_{i} ∈ CH^{∗}(V −S)/G. For this it seems natural to use the detailed information which we possess about BH since H is a product of the groups GL(n): namely, taking V to be the direct sum of copies of the standard representations of the various factors GL(n), we can arrange for (V − S)/H to be a product of Grassmannians, while making the codimension of S as large as we like. In particular, using such representations V of H, (V − S)/H is a smooth projective variety which has a cell decomposition, and the natural principal H-bundle over (V − S)/H is trivial over each cell. So also our H/G-bundle over (V − S)/H is trivial over each cell, and for each cell C ⊂ BH, (V − S)/G|_{C} is isomorphic to (H/G) × C and the Chow ring of (V − S)/G|_{C} is isomorphic to that of H/G. In particular, an element x ∈ CH^{∗}(V −S)/G which restricts to 0 in the Chow ring of a fixed fiber also restricts to 0 in the Chow ring of (V − S)/G|_{C}, where C is the big cell of (V − S)/H.

We now use the exact sequence

CH_{∗}Y → CH_{∗}X → CH_{∗}(X − Y) → 0

which exists for any algebraic variety X and closed subset Y. It follows that the above element x is equivalent in CH^{∗}(V − S)/G to a linear combination of Zariski closures of product varieties Z × C for cells C in (V − S)/H of codimension greater than 0. By induction on the dimension of these cells, and using that CH^{∗}(V −S)/G maps onto the Chow ring of every fiber H/G, we find that x can be written as a sum ∑ a_{i}x_{i} with a_{i} ∈ CH^{>0}(V −S)/H (representing the various cells of (V −S)/H of codimension > 0) and x_{i} ∈ CH^{∗}(V − S)/G. QED.

15. The Chow rings of the classical groups

For G = GL(n, C), then as we have explained in the proof of Proposition 14.2, BG can be approximated by smooth projective varieties (V − S)/G which have (algebraic) cell decompositions. For such varieties, it is well known that the Chow ring maps isomorphically to the cohomology ring [13], and hence we have CH^{∗}BG = H^{∗}(BG, Z) for such groups. In particular:

CH^{∗}BC^{∗} = Z[c_{1}]
CH^{∗}BGL(n, C) = Z[c_{1}, …, c_{n}]

Here and in what follows, we write c_{i} for the ith Chern class of the standard representation of a group G when this has an obvious meaning. These rings agree with MU^{∗}BG ⊗_{MU^{∗}} Z, for the trivial reason that MU^{∗}X ⊗_{MU^{∗}} Z = H^{∗}(X, Z) for all spaces X with torsion-free cohomology.

A much more interesting calculation, proved below, is that

CH^{∗}BO(n) = Z[c_{1}, …, c_{n}]/(2c_{i} = 0 for i odd).

This agrees with MU^{∗}BO(n) ⊗_{MU^{∗}} Z by Wilson’s calculation of MU^{∗}BO(n) [45]. It is simpler than the integral cohomology of BO(n), which has more 2-torsion related to Stiefel-Whitney classes. Let us mention that CH^{∗}BO(n) → H^{∗}(BO(n), Z) is injective, and in fact additively split injective. This is clear after tensoring with Z_{(p)} for an odd prime p, where both rings are polynomial rings on the Pontrjagin classes c_{2}, c_{4}, c_{6}, …, c_{2⌊n/2⌋}. To prove that the homomorphism

Z[c_{1}, …, c_{n}]/(2c_{i} = 0 for i odd) → H^{∗}(BO(n), Z)

is additively split injective 2-locally, it suffices to show that it is injective after tensoring with Z/2, given that the first groups have no 4-torsion. To prove the injectivity after tensoring with Z/2, we have

CH^{∗}BO(n) ⊗Z Z/2 = Z/2[c_{1}, c_{2}, …, c_{n}]
H^{∗}(BO(n), Z) ⊗Z Z/2 ↪ H^{∗}(BO(n), Z/2) = Z/2[w_{1}, w_{2}, …, w_{n}],

and c_{i} restricts to w_{i}^{2} [27], so the map is injective.

To make the above computation of the Chow ring of BO(n), we use Proposition 14.2, applied to the standard representation O(n) ⊂ GL(n). We find that the Chow ring of BO(n) is generated as a module over CH^{∗}BGL(n) = Z[c_{1}, …, c_{n}] by any elements of CH^{∗}BO(n) which map onto CH^{∗}GL(n)/O(n). But the variety GL(n)/O(n) is precisely the space of nondegenerate quadratic forms on C^{n}, and thus it is a Zariski open subset of affine space C^{n(n+1)/2}. By the fundamental exact sequence for Chow groups, the Chow groups of any Zariski open subset of affine space are 0 in codimension > 0 (and Z in codimension 0). It follows that CH^{∗}BGL(n) = Z[c_{1}, …, c_{n}] maps onto CH^{∗}BO(n).

It is easy to check that the relations 2c_{i} = 0 for i odd are true in CH^{∗}BO(n): this is because the representation O(n) → GL(n) is self-dual, and we have c_{i}(E^{∗}) = (−1)^{i}c_{i}(E) in CH^{∗}X for every algebraic vector bundle E on a variety X. (This is standard for algebraic geometers; for topologists, one might say that it follows from the corresponding equation in H^{∗}(X, Z) by considering the universal bundle on BGL(n) and using that CH^{∗}BGL(n) = H^{∗}(BGL(n), Z).)

To show that the resulting surjection

Z[c_{1}, …, c_{n}]/(2c_{i} = 0 for i odd) → CH^{∗}BO(n)

is an isomorphism, it suffices to observe that, by the calculation above, the composition Z[c_{1}, …, c_{n}]/(2c_{i} = 0 for i odd) → CH^{∗}BO(n) → H^{∗}(BO(n), Z) is injective, so the first map is also injective. This proves the description of CH^{∗}BO(n) stated above.

The symplectic group Sp(2n) is similar but simpler. As in the case of the orthogonal group, the quotient variety GL(2n)/Sp(2n) is the space of nondegenerate alternating forms on A^{2n} and hence is an open subspace of affine space. Thus GL(2n)/Sp(2n) has trivial Chow groups, and by Proposition 14.2, the homomorphism

CH^{∗}BGL(2n) → CH^{∗}BSp(2n)

is surjective. Thus the Chow ring of BSp(2n) is generated by the Chern classes of the natural representation, c_{1}, …, c_{2n}.

Since the natural representation of Sp(2n) is self-dual, we have 2c_{i} = 0 for i odd in CH^{i}(BSp(2n)). In fact, we have c_{i} = 0 for i odd. To see this, it is enough to show that the Chow ring of BSp(2n) injects into the Chow ring of BT for a maximal torus T, since the latter ring is torsion-free. Here the classifying space of a maximal torus can be viewed as an iterated affine-space bundle over the classifying space of a Borel subgroup B, so it is equivalent to show that the Chow ring of BSp(2n) injects into the Chow ring of BB. But BB is a bundle over BSp(2n) with fibers the smooth projective variety Sp(2n)/B. Moreover, the group Sp(2n) (unlike the orthogonal group) is special [8], so that this bundle is Zariskilocally trivial. Hence, taking the closure of a section of this bundle over a Zariski open subset, there is an element α ∈ CH^{r}BB, where r = dim Sp(2n)/B, such that f_{∗}α = 1 ∈ CH^{0}BSp(2n). This implies that f^{∗} : CH^{∗}BSp(2n) → CH^{∗}BB is injective as we want, since

f_{∗}(α · f^{∗}x) = x

for all x ∈ CH^{∗}BSp(2n).

Thus the Chow ring of BSp(2n) is generated by the even Chern classes c_{2}, c_{4}, …, c_{2n} of the standard representation. Using, for example, the map from the Chow ring to cohomology, which is the polynomial ring on these classes, we deduce that

CH^{∗}BSp(2n) = H^{∗}(BSp(2n), Z) = Z[c_{2}, c_{4}, …, c_{2n}].

16. Other connected groups

It is easy to compute the Chow ring of BSO(2n+1) using the previous section’s computation for BO(2n + 1). Indeed, there is an isomorphism of groups

O(2n + 1) ≅ Z/2 × SO(2n + 1),

where the Z/2 is generated by −1 ∈ O(2n + 1). Since BZ/2 can be approximated by linear varieties in the sense of [41], we have

CH^{∗}B(G × Z/2) = CH^{∗}BG ⊗Z CH^{∗}BZ/2,

where CH^{∗}BZ/2 = Z[x]/(2x = 0), x ∈ CH^{1}. Applying this to G = SO(2n + 1) and using the previous section’s computation of CH^{∗}BO(2n + 1), we find that

CH^{∗}BSO(2n + 1) = Z[c_{2}, c_{3}, …, c_{2n+1}]/(2c_{i} = 0 for i odd).

Since we also have MU^{∗}B(G × Z/2) = MU^{∗}BG ⊗_{MU^{∗}} BZ/2 for all compact Lie groups G, Wilson’s calculation of MU^{∗}BO(2n + 1) determines the calculation of MU^{∗}BSO(2n+1), and we find that CH^{∗}BSO(2n+1) agrees with MU^{∗}BSO(2n+ 1) ⊗_{MU^{∗}} Z. Pandharipande has another computation of the Chow rings of BO(n) and BSO(2n + 1), and he used these calculations to compute the Chow ring of the variety of rational normal curves in P^{n} [31].

The groups SO(2n) are more complicated, in that the Chow ring is not generated by the Chern classes of the standard representation. One also has, at least, the nth Chern class of the representation of SO(2n) whose highest weight is twice that of one of the two spin representations of Spin(2n): this class is linearly independent of the space of polynomials in the Chern classes of the standard representation even in rational cohomology. One can guess that the Chow ring of SO(2n) is generated by the Chern classes of the standard representation together with this one further class; Pandharipande has proved this in the case of SO(4) [31]. His calculation agrees with MU^{∗}BSO(4) ⊗_{MU^{∗}} Z, where MU^{∗}BSO(4) was computed by Kono and Yagita [21].

The group P GL(3) is even more complicated, in that MU^{∗}BP GL(3) is concentrated in even degrees but not generated by Chern classes, by Kono and Yagita [21]. So my conjecture predicts that CH^{∗}BP GL(3) is also not generated by Chern classes (as happened in the case of the finite group S_{6}; see section 4).

Finally, we can begin to compute the Chow ring of the exceptional group G_{2}. We use that G_{2} is the group of automorphisms of a general skew-symmetric cubic form on C^{7} [14], p. 357. Equivalently, GL(7)/G_{2} is a Zariski open subset of Λ^{3}(C^{7}) = C^{35}, and so it has trivial Chow groups. By Proposition 14.2, it follows that the Chow ring of BG_{2} is generated by the Chern classes c_{1}, …, c_{7} of the standard representation G_{2} ↪ GL(7).

References

[1] M. Atiyah. Characters and cohomology of finite groups. Publ. Math. IHES 9 (1961), 23-64.

[2] M. Atiyah and F. Hirzebruch. Riemann-Roch theorems for differentiable manifolds. Bull. AMS 65 (1959), 276-281.

[3] M. Atiyah and F. Hirzebruch. Vector bundles and homogeneous spaces. Differential Geometry, Proc. Symp. Pure Math. 3, 7-38. Providence: American Mathematical Society (1961).

[4] M. Atiyah and F. Hirzebruch. Analytic cycles on complex manifolds. Topology 1 (1962), 25-45.

[5] M. Atiyah and G. Segal. Equivariant K-theory and completion. J. Diff. Geom. 3 (1969), 1-18.

[6] F. Bogomolov. The Brauer group of quotient spaces by linear group actions. Math. USSR Izv. 30 (1988), 455-485.

[7] K. Brown. Cohomology of Groups. New York: Springer-Verlag (1982).

[8] Anneaux de Chow et applications, Séminaire C. Chevalley, 2e année. Paris: Secrétariat mathématique (1958).

[9] J.-L. Colliot-Thélène. Cycles algébriques de torsion et K-théorie algébrique. Arithmetic Algebraic Geometry (Trento, 1991), Lecture Notes in Mathematics 1553, 1-49. Berlin: Springer-Verlag (1993).

[10] P. Conner, E. Floyd. The Relation of Cobordism to K-Theories, Lecture Notes in Mathematics 28. Berlin: Springer-Verlag (1966).

[11] D. Edidin and W. Graham. Equivariant intersection theory. Preprint.

[12] H. Esnault, B. Kahn, M. Levine, and E. Viehweg. The Arason invariant and mod 2 algebraic cycles. J. Amer. Math. Soc., to appear.

[13] W. Fulton. Intersection Theory. Berlin, New York: Springer-Verlag (1984).

[14] W. Fulton and J. Harris. Representation Theory. New York: Springer-Verlag (1991).

[15] W. Fulton, R. MacPherson, F. Sottile, and B. Sturmfels. Intersection theory on spherical varieties. J. Alg. Geo. 4 (1995), 181-193.

[16] A. Grothendieck. Classes de Chern et représentations linéaires des groupes discrets. Dix Exposés sur la Cohomologie des Schémas, 215-305. Amsterdam: North-Holland (1968).

[17] A. Grothendieck. Problèmes ouverts en théorie des intersections. Théorie des intersection et théorème de Riemann-Roch, SGA 6, Lecture Notes in Mathematics 225, Exposé XIV, 667-689. Berlin: Springer-Verlag (1971).

[18] R. Hartshorne. Equivalence relations on algebraic cycles and subvarieties of small codimension. Algebraic Geometry, Proc. Symp. Pure Math. 29, 129-164. Providence: American Mathematical Society (1975).

[19] M. Hopkins, N. Kuhn, and D. Ravenel. Morava K-theories of classifying spaces and generalized characters for finite groups. Algebraic Topology (San Feliu de Guixols, 1990), Lecture Notes in Mathematics 1509, 186-209. Berlin: Springer-Verlag (1992).

[20] J.-P. Jouanolou. Une séquence exacte de Mayer-Vietoris en K-théorie algébrique. Algebraic K-Theory I: Higher K-Theories, Lecture Notes in Mathematics 341, 293-316. Berlin, New York: Springer-Verlag (1973).

[21] A. Kono and N. Yagita. Brown-Peterson and ordinary cohomology theories of classifying spaces for compact Lie groups. Trans. AMS 339 (1993), 781-798.

[22] I. Kriz. Morava K-theory of classifying spaces: some calculations. Topology 36 (1997), 1247-1273.

[23] P. Landweber. Coherence, flatness, and cobordism of classifying spaces. Proc. Advanced Study Inst. on Algebraic Topology (Aarhus, 1970), Vol. II, 256-259. Aarhus: Mat. Inst., Aarhus University (1970).

[24] P. Landweber. Elements of infinite filtration in complex cobordism. Math. Scand. 30 (1972), 223-226.

[25] I. Madsen and J. Milgram. Classifying Spaces for Surgery and Cobordism of Manifolds. Princeton: Princeton University Press (1979).

[26] A. S. Merkurjev. Comparison of equivariant and ordinary K-theories of algebraic varieties. Preprint.

[27] J. Milnor and J. Stasheff. Characteristic Classes. Princeton: Princeton University Press (1974).

[28] D. Mumford, J. Fogarty, and F. Kirwan. Geometric Invariant Theory. New York: Springer-Verlag (1994).

[29] M. Nakaoka. Cohomology theory of a complex with a transformation of prime period and its applications. J. Inst. Polytech. Osaka City University 7 (1956), 51-102.

[30] M. Nori. Algebraic cycles and Hodge-theoretic connectivity. Invent. Math. 111 (1993), 349-373.

[31] R. Pandharipande. The equivariant Chow rings of O(k), SO(2k + 1), and SO(4). To appear in J. Reine Angew. Math.

[32] K. Paranjape. Cohomological and cycle-theoretic connectivity. Ann. Math. 139 (1994), 641-660.

[33] D. Ravenel, S. Wilson, and N. Yagita. Brown-Peterson cohomology from Morava K-theory. Preprint.

[34] J.-P. Serre. Sur la topologie des variétés algébriques en caractéristique p. Symposium internacional de topologia algebraica, Mexico (1958), 24-53; Oeuvres, v. 1, 501-530.

[35] J.-P. Serre. Linear Representations of Finite Groups. New York: Springer-Verlag (1977).

[36] R. Stong. Notes on Cobordism Theory. Princeton University Press, Princeton (1968).

[37] A. Suslin. Algebraic K-theory of fields. Proceedings of the International Congress of Mathematicians, Berkeley, 1986. Providence: American Mathematical Society (1987), 222-244.

[38] M. Tanabe. On Morava K-theories of Chevalley groups. Am. J. Math. 117 (1995), 263-278.

[39] M. Tezuka and N. Yagita. Cohomology of finite groups and Brown-Peterson cohomology. Algebraic Topology (Arcata, 1986), Lecture Notes in Mathematics 1370, 396-408. Berlin: Springer-Verlag (1989).

[40] C. B. Thomas. Filtrations on the representation ring of a finite group. Proceedings of the Northwestern Homotopy Theory Conference (Evanston, 1982), Contemp. Math. 19, 397-405. Providence: American Mathematical Society (1983).

[41] B. Totaro. Chow groups of linear varieties. To appear in J. Alg. Geo.

[42] B. Totaro. Torsion algebraic cycles and complex cobordism. J. AMS 10 (1997), 467-493.

[43] W. Waterhouse. Introduction to Affine Group Schemes. New York: Springer-Verlag (1979).

[44] W. S. Wilson. The Ω-spectrum for Brown-Peterson cohomology, part I. Comment. Math. Helv. 48 (1973), 45-55.

[45] W. S. Wilson. The complex cobordism of BO(n). J. LMS 29 (1984), 352-366.

Department of Mathematics, University of Chicago, 5734 S. University Ave., Chicago, IL 60637

    E-mail address: totaro@math.uchicago.edu